A Complete Bibliography of Publications in Linear Algebra and its Applications: 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

11 April 2024 Version 2.04

Title word cross-reference

(AB)^{+} = B_{mr}^{−}A_{lr}^{−} [WHG79]. (A × I + I × B)x = g [LNPU73]. (Ay = λBy) [TW72]. (k) [Cha79]. (k, n) [MT79]. (n − 1) [Kap72]. (v, q + 1, 1) [Qua75]. −1 − \sqrt{2} [Hof77]. 0 [HP78, JGK79]. 1 [Har75b, HP78]. 2 [Har75b, Sto79]. 2 × 2 [Est79, Tau75, Win71a]. 3 × 3 [AYP79, BC77]. _{2} [Coh78]. _{n} [Cou73]. A [Nic79]. A^{2} = dI + λJ [Lam75]. A^{2} = J [Rys70]. AB + BA [Nic79]. AH + HA^{∗} = A^{∗}H + HA = I [GG77]. AX − Y B = C [BK79a]. AX = B [Por76, PM79]. AXC = B [PM79]. B [Nic79]. BA^{+}C [Har75b]. C [GS79]. C^{2} [Fer78a]. D [Dat78c]. ϵ [Bre74, Bre75]. G [CV73a, CV73b, CV74]. γ [Mey70]. GF(q) [Qua75]. GF(q)[x]/(a(x)) [Cla78a]. H [Pan79b]. h(T)(x) [LV75, LV76]. k [Bea70, Chr79, GS76, Smi77]. L [Rya79, Tho70]. Λ [BK79b, Mar79c]. λ = 1 [Doo75]. LU [Cry73]. M [Mar79b, NP79, Rot79a, Smi79, wK77a, wK77b, MS78, Ple76, Ple77, Rag78, Wat72, Wil77, BM76, JGK79]. M = (AD; BC) [HM75]. n [Dav79, Kap72, MR73, Pin79]. n × n [BCL75]. ω [Var77]. P [GZ79a, MR73, Hed75, HW79b, Laf75]. P^{∗}KP + L [Ber74]. Q [KW79].

Q(h(T)(x)) [LV75, LV76]. Q(x) [LV75, LV76]. qd [Hou71b]. QR [DT71, Gen75]. QZ [Wil79]. r [Kim75]. R^{n} [O’B75, Uhr77]. S [MR73]. SA + A^{∗}S = S^{∗}B^{∗}BS [CD79]. T A − BT = C [Hea77]. τ [Var77]. T B + CT = −A [HW79c]. XA + AY = F [HMW79]. Z [Pan79a]. ||A^{−1}||_{∞} [Var76a].

-algorithm [Bre74, Bre75]. -BIBD [Qua75]. -bundles [Coh78]. -characters [Cha79]. -circulant [Smi77]. -dimensional [MR73]. -factorization [Cry73]. -functions [Rya79, CV73a, CV73b, CV74, MR73]. -gons [Dav79]. -inverses [Mey71]. -matrices [KW79, Mar79b, Mar79c, NP79, Pan79b, Rot79a, Smi79, Var77, GZ79a, wK77a, wK77b, MS78, Pan79a, Ple76, Ple77, Rag78]. -matrix [But72, Ple77, Wat72, Wil77]. -measure [Chr79]. -Minors [Hed75]. -numerical [GS76, GS79]. -orthogonal [BK79b]. -potent [Kim75]. -Square [Kap72]. -stability [Dat78c]. -symmetric [JGK79]. -Widths [Pin79].

1977 [Ano77i]. 1978 [Ano78k].

A. [Ste79c]. Abschätzung [Sch78]. absolute [Fri79, Par79, Tho71]. Accurate [Mat75]. achieves [Hed75]. actions [Hof76b]. Addendum [Ano74a]. addition [MT75]. Additive [Tau75]. ADI [LNPU73]. adjoint [Djo73]. adjoints [Hal71]. affine [Hof76b, Put79]. afforded [IM79]. Alg [Lon79]. algebra [Ano79k, Hol79, Laf79, Lyl78, Mer78]. Algebraic [BS75, Bri78, Rot75, Els79, Lyl78, Win71b]. algebras [Djo75, Hir74, Ion76, Kah71, LC70, MM72a, O’R77, Rob75]. Algebren [Hir74]. Algorithm [DT71, Bal71, Bre74, Bre75, CGG77, DD76, GM73, Gre73, Hou71b, Hu70, Wil79]. algorithmes [Mor76]. algorithms [DD75, Els79, Mor76]. Almost [Pad76, Ruh75]. ALPS [BW78]. Alston [Ano77a, Ano78a, Wil75a]. alternative [Mey78b, Mih71]. Among [Par76]. analogue [Bar75]. Analysis [Gen75, GLR78a, GLR78b, DR76, Sty73]. Analytic [HP79b]. angles [CP79]. anisotropic [Hea75]. Announcement [Ano77b]. Antiderivations [MM72a]. antisymmetrien [Bod79]. antisymmetries [Bod79]. application [BK73, Kru77, Ros76a, Sil70, Usm71, Wii78]. Applications [Bre76, Lan77a, Lan77b, Mer78, Van71, And79a, BBI73, Bri78, DD76, Fan75, FL74b, Neu78]. approach [Dav77, Ger77, Hoc70, Naz77, Par70, Pul71]. approximate [Kov76]. approximation [Sta75]. Arbitrary [Hua78, Sch79, Deu71, Fre73, Joh71a]. Archimedian [Bye70]. Area [Smi71]. arguments [Var76a]. arising [Law75]. arithmetic [Kru77]. Arnold [Was77]. arrays [Kru77]. Artinian [McD78]. assignment [BL78, Cru79]. Associated [Var73, CF79, Mer71]. asymptotic [Hu70, Was77]. August [Ano77i]. Author [Ano70a, Ano71a, Ano73a, Ano73b, Ano75a, Ano75b, Ano75c, Ano76a, Ano76b, Ano76c, Ano77c, Ano77d, Ano77e, Ano78c, Ano78d, Ano78b, Ano78e, Ano78f, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f]. automatic [Ban79]. automorphism [Hor78b]. automorphisms [Est79, Gil77]. averaging [Fla74]. Award [Ano77a, Ano78a, Wil75a].

B [Lee75, SS73]. B-splines [Lee75, SS73]. Banach [Kov74, Kov75, Vid76]. Band [GT79, Met73, Usm71, YI79]. Bandmatrizen [Met73]. based [Fre73]. bases [Bra70]. basic [Gib76, Naz77]. be [Far77]. behavior [Cam79, Tho76b]. Best [JO75, LNPU73]. between [BK79b, Djo73, Djo75, FO78, Hor75, Par75, Sin79, Var73, Vit79]. BIBD [Doo75, Qua75]. bilinear [AS79, BD78, Laf75, Zel79b]. biquadratic [Cal73, Cho75b]. birth [HM70]. Block [Eis73, Eis76, Kos78, Uhl73, Bar76, Pow76, Sha79, Smi77]. block-partitioned [Bar76]. blocks [Hal76b, RR73]. blocs [RR73]. board [Ano72, Ano73d, Ano74b, Ano75d, Ano76d, Ano77f, Ano77g, Ano78g, Ano70b, Ano71b, Ano73c, Ano75e, Ano75f, Ano76e, Ano76f, Ano77h, Ano78h, Ano78i, Ano78j, Ano79g, Ano79h, Ano79i, Ano79j]. Boolean [FG77, Gra75, KR78, KR79a, RR75, Rob78a, SS76]. booléens [Rob78a]. bordered [Hal76b]. Bound [Sch70, Var75]. boundary [MF78, Nat73, Usm71]. Boundedness [Wil70]. bounding [Var76a]. Bounds [BG74b, BG76a, Fie72, FM78, Joh74b, Rob73, Ste79b, TW78, Wil75b, Bye70, Nic79, Nwo78, Pat76, vLvEB78, FL74a]. Bruijn [HP79a]. bundles [Coh78]. burst [SD77].

calculating [Els76b]. calculation [Tch79]. calculations [TW72]. Calculus [Dav73, Sil70]. Camion [Lev72a]. can [Heu73]. Canonical [Hol79, Uhl76, Van79, Hea73, Hoc70, Oko79, Was77, Wil79, GLR78a, Joh74a]. capacity [Ste77]. Cardinal [LS73, SS73, Lee75, LS73]. cas [RR73]. case [Joh74a, RR73]. Cauchy [Mar79a]. Cayley [Fan72]. central [Cat78, Weh77]. centralizer [Smi74, Smi76]. centrosymmetric [CB76, Cru77, PBA73]. certain [And79a, And73, Bar76, CPW74, Doo70, HW79a, Hil77, Joh74b, Kan78b, Pad76, Par75, PR77, Sin77, Sto79]. chain [Sta75]. chains [Har75a, Paz75]. change [BM79]. Chaotic [Don71, Rob76]. chaotiques [Rob76]. characteristic [Bax78, BG74b, BG76a, Ful73b, Ful79, KP74, Tho76c, Wii78, dO72, dO74]. Characterization [Cia70, Djo75, Ful73a, Cha75, Cru75, Doo75, Fri79, Pan79b, RS74, Rot79b, Uhr77]. characterizations [DC79, Els76a, Fun79, Kan78a, Ple77]. characterizing [Doo70]. characters [Cha79, Fre73]. Chebyshev [Uhr77]. Chipart [Dat78a]. circulant [Sea79, Smi77, Zel79a]. circulants [Cha73, Lam75, Wan79]. circular [CP79]. class [Aba79, AT76, And78, AHS76, Cia70, Fer78a, Mee76, MW79, Moh76, vK70]. classes [Cha79, CHL70, Ful73a, Ful73b, MR73, Mul79, Reh77, Rot79b, BW78]. classical [Hoc70]. Classification [Djo78, Rot79a]. classified [FMKL79]. Closeness [Ruh75]. codes [SD77]. coding [Lyl78]. Coefficients [Lan77a, Lan77b, Hol77, Hol78]. Cohn [Dat78b]. collections [CV74]. collineations [Cat78]. column [BC73, HP78]. columns [Mar79d]. Combinations [Maj79, Kap75]. combinatorial [Bri78, Cru77, Hof76a, Pad76, Par70]. Common [Bar77, Bar70, MB78]. commutative [BEK78, BG78, Chi77, Cou73, Est79, Gus79]. commutativity [Par75]. commutators [Max71, Rob71, Tau75]. commuting [Dav70, LNP71, Wat76]. compact [Gil77]. companion [Bar75]. Comparison [Sen75]. Compatibility [Kov76, Joh77b]. Compatible [Kov77]. Complement [Cot74, BJ78, Hay70, Wat72]. complementarity [Kan78a, Kan78b, Kos78, Mur72, Pan78]. complementary [Moh76, Mur72]. complements [And79b, Ost71]. Completely [Cho75a, Rag78]. Complex [KS78, Bar71, Cho75a, DR77, Heu73, KR73, Laf77, LC70, MN79, Sei74, Tho79b, Tho79c]. complexity [AS79, KR79a, Kru77, Win71b, vLvEB78]. Computation [Bre75, Van79, Laf75, Zel79b]. Computational [Law75]. computer [Tch79]. Computing [FG77, Hua78, DD76, HW79b, LNPU73, Ros76b]. Concavity [And79a]. concepts [Sch77]. concerning [Lev72b]. Condition [GJ70, Mar73, Sau77, FL74a]. Conditions [Dia79, Kwo77, dOdS79, Far77, Mar79c]. cone [Bar73a, BT73, Hor78b, Hor78a, Pan78]. cones [Bar73b, BL75, BF76, Bar78, BIL78, BBI73, HFP76, LS75, Mur72, O’B75]. congruence [Cai73, Joh74a, Ng76]. conjecture [KR79b, Lon71, Min79]. conjugacy [Roi78]. Connection [Var73, Els79]. considerations [Har77]. considered [Cla78a]. Consistency [vdSV79, Fan75]. consisting [Djo78]. Constant [Lan77a, Lan77b, FL74b, HP78, Sin77]. constrained [Mar79d, You79]. constraint [Wil70]. constraints [Far77]. Construction [GW76, Hea75, Hoc79, Ros76a, SS76]. constructive [Hoc70]. containing [MA79]. Context [San72]. Context-free [San72]. continuant [Hea70]. Continuity [CM75, Cam77]. continuous [Sta75]. Contraction [RR73, Rob76, Rob76]. control [FW74]. controllability [CH76, Wim74]. Convergence [Rob76, Naz75, Sto79]. Convex [BG76b, Eck75, Par70, Smi71]. convexity [AYP79, CF72, Fox78, Sau77]. copositive [CHL70]. correcting [SD77]. Corrections [Bre76]. criteria [Bal78b, CF72, Fer78a]. criterion [Dat78a, Laf79, Met73, Ng76]. Cross [Tam76, Tam75]. Cross-Positive [Tam76, Tam75]. crypto [Hol78]. crypto-unitary [Hol78]. cryptocontractions [Hol77]. curious [BK73]. curves [Hil77]. Cyclic [Dav79, ES73, FT79, Cha70, Fei77, Naz75, Ste77].

data [GW76, Hoc79, dBG78]. death [HM70]. decomposable [BH79, Dia79, Mer77, Min72, dOdS79]. Decomposition [Rad75, And74]. Decompositions [Gen75, Kru77]. defined [And74]. définies [Far76]. Definite [GV79, Ste79b, And79a, AY73, Bal70, Djo76, HW79b, Joh73, KP74, Nic79]. definiteness [AY75]. degrees [Car72]. Dénes [Gal78]. density [SD77]. dependence [Sil70, Ste77]. derivative [MSH77]. derivatives [DM78, Pao73a, Pao73b]. designs [Gra75]. detection [Met73]. determinant [AHS76, Bre76]. Determinants [Est79, Joh74b, Vit79]. determined [Web74]. Deterministic [Arn76]. diagonal [Ach78, BC77, BVW78, ES73, Neu79, PR77, Tho79c, TW78, Var76a, Var76b]. Diagonalization [Uhl73, Eis73, Eis76, Sha79]. diagonals [Pan79b, Sin77]. Dieudonné [Bre76]. difference [AG75, Ber75, Ber76, Cam79, CF79, FM78, Ros76a, Tur77, Lan77b]. differences [Vit79]. Differential [Lan77a, FL74b, Nat73, Was77]. dilation [Dav77]. dimension [BC75, Eck75, Gro77]. dimensional [AS73, Bar73a, Fei77, FO78, Moo74, MR73, Van76, Vid76]. dimensions [AH79, Ben70, Lon79]. Diophantine [CP79]. Direct [Oko79, O’R77, Par79]. Dirichlet [Olk79, Rya79]. Dirichlet-type [Olk79]. disks [JO75]. Dissipative [Tho75a, Fan72]. distance [FMKL79]. Distortion [Hol77, Hol78]. distribution [Arn76]. d’itérations [Rob76]. divisible [Far76]. divisibles [Far76]. division [DM79]. divisor [Bar70, MB78]. Divisors [GLR78a, GLR78b, Car72]. Domains [MA79, Zen77, vdS79]. dominance [Joh79, Neu79, Var76a, Var76b]. double [Max71]. Doubly [Ach78, Sin77, BG75, BW79, BG76b, Chr79, Fie72, Gib76, Gib78, Hed75, Joh71a, Joh74a, Lev72b]. Drazin [CM75, Cam77, CM78, Gre73, Ros76b]. dual [Aba79, BF76, BIL78, GP79, Gra75]. duality [Bal71, Bri78]. Dulmage [Kim79]. D’Yakanov [Ban79]. dynamics [Ros76a].

edges [Hof72]. Editorial [Ano72, Ano73d, Ano74b, Ano75d, Ano76d, Ano77f, Ano77g, Ano78g, Ano70b, Ano71b, Ano73c, Ano75e, Ano75f, Ano76e, Ano76f, Ano77h, Ano78h, Ano78i, Ano78j, Ano79g, Ano79h, Ano79i, Ano79j]. effective [Bal78b, Ng76]. eigenelements [Bre75]. eigenproblem [Wer70]. eigenspaces [Rot75]. Eigenvalue [Fri75, Fri77, Hil77, Nic79, Ste79b, Els79, Hal76a, Hof77, MN77]. Eigenvalues [CB76, Hof72, Pai74, Ruh75, Van71, Var73, Arn76, Ber75, Ber74, Doo70, Eis76, Fie72, FM79, HM70, KS78, Loe71, Mor76, Nwo78, Sch78, Sin79, Smi71, TF71, Tho76b, Ves79, Zen77, vK70, JO75]. eigenvector [Wer73]. Eigenvectors [And73, Hal71, CB76, FL74b, Sen75]. eigenvektors [Wer73]. eigenwerte [Sch78]. Eight [BP76]. Einschließung [Wer73]. electrical [AT76]. Elementary [GS77, Ber74, Car72, Dat78a, DB71, Joh77a, Kap72, vLvEB78]. elemente [Hir74]. Elements [Par76, Bal78a, HK76, Hir74, Mar79d, Smi74, Tho79c, TW78, Vit79, vK70, Joh71a]. Elimination [Coh74]. embeddings [DC79]. Enclosure [Wer73]. endomorphisms [MR76, Rob71]. Eneström [ASV79]. entire [MOV78]. entries [BVW78, IM79]. Enumeration [BC73, CH77, Ful73a]. EP [Bal75, Duk70, HK76, Mey70]. Epimorphisms [Pra76]. EPr [Kat75]. equal [Ach78, Nic75, Wes72]. equality [Bal78a, Dia79, dOdS79]. Equation [Por76, BK79a, CD79, Hal79, HMW79, HW79c, Rys70, Spa78]. Equations [Kov77, Lan77a, Lan77b, Van77, Vet75, Bar77, Ber75, Ber76, BBI73, Cam79, Chi77, Dau78, FL74b, FM78, GG77, GZ79b, Har78, Kov76, Nat73, OS72, PT79b, Rob76, Ros76a, Sha77, SW73, Was77, Rob76]. equimodular [Eng73]. Equivalence [Mul79, Ger77, GZ79b, Joh74a]. Erdos [HP79a, KR79b]. Error [Gen75, Rob73]. errors [SD77]. Essentially [Joh71a, Joh74a, Joh77a, Kel71]. estimate [Sch78]. Estimates [Coo75, Nwo78, Mat75]. Euclidean [BF76]. evaluation [BD78]. exact [FL74b]. examples [Sto79]. exceeds [Hof77]. existence [Bax78, HR74, Mar79d, Ptá76, TW72]. Expansion [Lan71]. Expansions [Tan75]. Explicit [Dav73, WHG79]. exponent [Cou73]. expression [Mey78b]. extension [GR79, GK76, Kim79]. Extensions [FO78, OP79, Uhl79, Dav76, Olk79]. exterior [MM72a, MR76, MA77]. Extrema [Ber74]. Extremal [BC75, Dav76, Mar75]. Extreme [HFP76, O’B75].

F [Min79]. faces [Bar73a]. facially [BIL78]. Factorization [LM78, LV75, LV76, Smi79, Wii78, Bos75, Cry73]. factorizations [Joh77a, wK77b, Naz77]. Factorizing [Fle76]. factors [Tho79a, Uhl74]. faible [RR73]. families [GK76]. family [AS79]. faster [HW79b]. feasibility [HP79b]. Fiedler [RS74]. field [BG75, Bos75, BCL75, Bra79, Ful73a, Gib76, LV75, LV76, MA79, Siz77]. fields [BM76, Cor75, Ful73b, Ful79, Joh71a, Mul79]. Final [Ano77a, Rot79b]. Finding [Van71]. Finite [Weh77, AS73, Bar73a, BCL75, BG78, Bra79, Cor75, Ful73a, LV75, LV76, Lon79, Moo74, Mul79, Sta75, Uhr77, Van76, Vid76]. finite-dimensional [AS73, Vid76]. first [Mey78b]. five [Usm71]. five-band [Usm71]. fixe [Rob76]. fixed [Rob76]. fluid [Ros76a]. Form [GLR78b, Uhl76, Van79, Dav76, Far77, GM73, Hea73, Hoc70, Wii78, Wil79]. formal [Hir74]. formal-real [Hir74]. formal-reellen [Hir74]. Forms [GLR78a, Uhl79, And74, AS79, BD78, Cal73, Cho75b, Coh78, Dav79, DR77, Fox78, Hor75, Joh74a, Laf75, Zel79b]. formula [Fla74, Ost71, Tho76a]. formulas [FMKL79]. Foster [Fla74]. four [Doo70]. Frame [Gre73]. free [Pra76, San72]. French [Far76, Mor76, RR73, Rob76, Rob78a]. Fritz [KR73]. Frobenius [BS75, Hou71b, RR73, Rob78a, Sch77]. Fuglede [DGS76]. Fujiwara [Dat78b]. Fulkerson [Cru75, HP79b]. full [Sch77]. function [Kap72, PJ75]. Functional [Dav73, OS72]. functionals [Sto79]. Functions [Hou71a, Par76, Sta78, AT76, BCL75, Bra79, CP79, CV73a, CV73b, CV74, CF72, DB71, Dow71, Eng73, Fer78a, Fre73, Heu73, Hil77, IM79, LS73, LC70, LNP70, MM72b, Mer75, Mer78, MW79, MOV78, MR73, Paz75, Rya79, SS73]. Fundamental [Lan77a, Lan77b]. Further [SS79, Pao73a].

Galois [Est79]. game [TW72]. games [Rag78, Wer70]. Gatlinburg [Ano77b]. Gaussian [Coh74]. Gel’fand [Hal79]. general [DM79, Els79, GJ70, Tau72]. generalization [BG75, BG72, Lev72a, MG70, Mer71, Reb74, RS71, Rys70]. Generalizations [BG74a, Car76, Wim75, HW79c, Neu79, Wan79]. Generalized [And79b, CH76, CPW74, DB71, Fan72, Flo79, Ful79, Hea73, NP79, PT79b, Ste79b, Tan75, Van77, Van71, Wal75, AYP79, Bat78, Cam77, Cla78b, CF79, Eng73, FG77, Fre73, GS77, Hal76b, HR74, Hea75, Joh77b, Joh79, KP74, Law75, MB78, Mey70, Mih71, MP79, Mül73, Nat73, Ong76, RR75, Wat72, Wer79]. generalizing [Har75a]. Generate [Uhl76]. generated [Ara79]. generators [Ion76]. genetic [Hol79]. Geometric [Dav77, Hil77, Pul71]. geometry [KW79]. German [Bod79, Els74, Hir74, Met73, Mül73, Sch78, Wer73]. Gerschgorin [Ber75, Ber76, Joh71b, JO75]. Gersgorin [Pow76]. Gershgorin [vdS79]. GF [Hof76b]. given [Ben70, CH77, Hoc79]. Givens [Gen75]. Global [How78]. gons [Dav79]. Grade [Sha77]. grammars [San72]. graph [Doo75, Hof72, KS78, Ste77]. Graphs [Ple72, Doo70, DC79, Hof77]. Grassmann [Kah71]. Greatest [Bar70, MB78, BG74b, BG76a]. Group [HW79a, DD75, DM79, Hor78b, KR79a, MG70, Vid76, Wii78]. group-theoretic [KR79a]. groupe [Tho75c]. groups [Dix76, Hof76b, MN70, Weh77, Wii78]. Gunn [Ban79].

H [Cai73]. H-stable [Cai73]. Hadamard [And79a, Ara79, Sty73]. halbgeordneten [Els74]. harmonic [DR76]. Haynsworth [Ost71]. height [Weh77]. Hermite [Lee75, Leh71, LS73, SS73]. Hermitian [AY75, Cai73, DT71, Djo75, Far76, Fox78, Hil73, Hor75, Kwo77, Maj79, Mar75, Pai74, Par79, PT79a, Por75, Sch79, TF71, Tho76c, vK70]. Hermitiennes [Far76]. Hessenberg [DD76, Hua78, Ike79, Par73]. Hidden [Pan79a]. High [HMW79]. High-order [HMW79]. higher [MF78, Rob71]. higher-order [Rob71]. Hilbert [Sto79]. Hill [Hoc75]. Hodge [MR75]. Hoffman [Lev72a]. Homogeneous [BIL78, Far77]. Householder [Ano77a, Ano78a, Ste79c, Wil75a]. hull [Smi71]. Hurwitz [Dat78b]. hybrid [MT75]. hyperbolic [Tur77]. hyperinvariant [FHL77]. Hyperpower [Tan75].

ideal [McD78]. ideals [Hor78a]. idempotent [Bal78c, JGK79, Kim74]. idempotents [But72]. Identities [Hou71b]. II [Ber76, BG76a, BW79, CV73b, GLR78b, Joh74a, Kat75, Kov75, Lan77b, LM78, LV75, dOdS79]. III [LS73, Mar73, Ong76]. Imbedding [Mar79c]. inaccurate [Kov76]. Inclusion [GS76, GS77, Jen73, Zen77]. incompatible [LNPU73]. indecomposability [Sch77]. Indecomposable [LS75, Tam78]. indéfiniment [Far76]. indefinite [Fle76]. independence [Hal76b]. independent [NS79]. Index [Ano77d, Ano70a, Ano70c, Ano71a, Ano71c, Ano73a, Ano73b, Ano73f, Ano73g, Ano75a, Ano75b, Ano75c, Ano76a, Ano76b, Ano76c, Ano77c, Ano77e, Ano78c, Ano78d, Ano78b, Ano78e, Ano78f, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano79m, Rot79a]. induced [OS72]. induction [Ptá76]. Inequalities [Car72, De 71, Fre73, Rya79, AY73, Els74, Fan75, Hor75, Joh79, Mar73, PT79a, Tho72, Tho75b, Tho79a]. inequality [Joh73, MM72b, Mar79a, Sei74]. Inertia [Wim74, BC77, CH76, JB70].

Infima [Var73]. infinite [FO78, SW73]. infinite-dimensional [FO78]. infinitely [Far76]. infinitely-divisible [Far76]. infinitesimal [Rob75]. inner [Win71b]. Inners [BJ78]. integer [Bal71, FT79, GP79, Hu70]. integers [BM76]. Integral [Bat78, Kwo77, Pad76]. integrality [GP79]. integrals [Olk79]. interior [BBI73]. Interlacing [Tho72, Tho79a]. interpolation [FM78, Lee75, LS73, SS73, Tho76a]. intersection [Cun77]. interval [CGG77, RBI70]. Intrinsic [LC70]. invariance [Bea70, Har75b, Ong76, RW77]. invariant [HM78, Kap72, Tho79a, Web74]. invariants [Coh78]. Inverse [Els76b, Fri75, Fri77, Hal76a, CH75, Coo75, FG77, Fu75, Hoc75, HM75, Jen73, JLR79, Lan77c, Law75, LNP71, MSH77, MS78, Mih71, Mor76, NP79, Nwo78, Ros76b, Smi78, Wil77]. inverse-positive [JLR79]. inverse-positivity [NP79]. Inverses [Ike79, KR78, Tan75, Bar79, Bat78, Cam77, CM78, CPW74, CF79, Deu71, Flo79, Ful79, GT79, Hal76b, HR74, Har78, HW79a, wK77a, LM79, Mee76, Mey70, Mey71, MP79, PT79b, RR75, Smi77, Wal75, Har75b]. Inversion [YI79, Bar76, Dow71, FMKL79, Wer79]. Invertibility [Lan73]. inverting [Sea79]. involutions [Ful73a, Ful73b, GHR76, Kov79, Rad75, Wii78]. Involutory [BG78, Ell79]. involving [CP79, GS76, Mer75, PT79a]. irreducibility [Cha75, Chr79, Els76a, Sch77, Van76]. irreducible [BG72, BG74b, BG76a, Els76b, Hea70, LV75, LV76, Sen75, Wer73]. irreduziblen [Wer73]. Isolated [Dau78, JO75]. isometries [Rob75, Vid76]. isotropic [Pal72]. iterated [Wil75b]. iteration [Ban79, Els76b, Jen73, MN77, SS76, Spa78]. iterations [OP79, Rob76]. Iterative [Tan75, HMW79, Neu78, Ptá76]. IV [Bal70, BG76b]. IX [Tho72].

Jacobi [FM79, Fu75, GW76, Hal76a, Hoc79, Naz75, dBG78]. Jacobians [OS72]. John [KR73]. Jordan [AH79, Hea73, Hir74, Wii78, Wim78]. Jordan-Algebren [Hir74]. July [Ano74a]. June [Ano78k].

Kakeya [ASV79]. Ketten [Wim78]. knots [FM78]. König [Gal78, Sch77]. Kreiss [GK76]. Kriterium [Met73]. Kronecker [AM76, Ara79, Van79, Wil79].

L [BVW78, RS71]. lacunary [Deu73]. Lambda [Lan77a, Lan77b]. Lambda-Matrices [Lan77a, Lan77b]. Large [Van71]. Lat [Lon79]. lattice [Bar73a, FHL77]. Laurent [Lan71]. law [Arn76, BG74a, SS79, WHG79]. leading [Ple76]. least [Hof77, LNPU73]. least-squares [LNPU73]. leave [Kap72]. lemma [AM76]. length [BK78]. Leontief [Sie79]. Letters [Ano79k]. Levitan [Hal79]. Lie [Rob75, Ion76]. Liénard [Dat78a]. life [Gal78]. like [PJ75]. Limit [Cam79]. line [Doo75]. linéaires [Rob76]. Linear [Bea70, BBI73, Bot78, Chi77, Ful73b, GV79, Hil73, Kan78a, Kov77, LO78, Maj79, Ong76, RW77, SW73, Vet75, Wat76, Ano79k, AS73, Ber75, CGG77, Cho75a, Cla78a, DF75, DD75, Deu71, Dix76, Djo73, Djo78, DM79, Duk70, DR76, Eav73, Fan75, Far77, FHL77, FO78, Ful73a, Heu73, Hoc70, Kan78b, Kap75, Kim74, Kim75, Kov76, Kov79, LV75, LV76, Lyl78, Man79, Moh76, Neu78, Pan78, RBI70, Rob73, SD77, Sha77, Sil70, Smi76, Ste77, Tch79, Van76, Weh77, Wes72, Wil70, Wim74]. linear-algebra [Lyl78]. linear-programming [CGG77]. linearly [NS79]. linked [Hol79]. Liouville [Olk79]. List [Ano73e, Ano76g, Ano77i, Ano78k, Ano79l]. Local [TC78, Dix76, Ger77, Joh76a]. localization [Kov74, Kov75]. Location [Bar71, Cla78b, Hil77, Loe71]. loci [Hol79]. Logarithmic [Pao73b, DM78, Pao73a]. low [SD77]. low-density [SD77]. Lower [Pat76, Var75, vLvEB78]. Lyapunov [Bar77, CL74, CD79, Joh76a, Loe76, Wim75].

M. [RS74]. Manifestations [Cot74]. Manifolds [Kov79]. map [Djo78]. mapping [Tch79]. mappings [Kim75, MR73, Tur79]. maps [And79a, BT73, Bot78, Cho75a, Djo75, Hor78a, RW77, Wat76]. Markoff [Coh78]. Markov [Har75a, Paz75, Sch77, Sta75]. Mathematical [Har77, Ptá76]. Matric [Por76]. Matrices [DT71, Fri75, Hua78, Ike79, Kel71, Lan77a, Lan77b, Maj79, Moy77, Nic75, Pai74, Par76, Pin79, Ruh75, Tam76, Uhl73, Uhl76, Van71, Ach78, Afr74, And78, And79a, AHS76, And73, Ara79, AY75, AYP79, BC77, Bal78a, Bal70, Bal75, Bal78c, BL75, Bar76, Bar79, Bat78, Bax78, Bea70, BG75, Ben70, BW78, Bot78, Bra70, BC73, BCL75, BG78, BW79, BC75, BG76b, BH79, BK73, Cai73, CB76, CH77, Cha70, Cho75a, Chr79, Cia70, CPW74, CF79, CHL70, Cru75, Cru77, Cry73, Cry76, DB71, De 71, DGS76, Deu78, Djo76, DR77, Ell79, Els76a, Eng73, Est79, Fan72, Far76, Fie72, Fle76, FM79, FMKL79, Fu75, Ful79, GZ79a, Gib76, Gib78, GTZ75, GK76, Gre74, GT79, Hal76a, Hal76b, HW79a, Hei78, Hil73, Hoc75, HM70, HM78]. matrices [HW79b, HP78, JGK79, Joh71a, Joh74a, Joh77a, Joh73, Joh74c, JLR79, Joh71b, Kan78a, Kap72, Kap75, KP74, Kat75, KW79, KR78, KR79a, KDW74, eK73, wK77a, wK77b, Kwo77, Laf77, Laf79, LM78, Law75, Lev72b, LO78, Loe71, LM79, LS76, LNP70, LNP71, Man74, MT79, Mar75, Mar70a, Mar70b, Mar79b, MB78, Mar79d, Mar79c, MN79, McD78, Mee76, Met73, Mey70, Mey71, MS78, MW79, Mih71, Min72, MT75, MP79, Mul79, Neu76, NP79, Neu79, Nic79, O’B75, Oko79, Ong76, Ost74, Pan79a, Pan79b, Par73, PT79a, Ple72, Ple76, Ple77, PB79, Ptá79, Pul71, PR77, RW77, Rad75, Rag78, Reh77, RS74, Rob78b, Roi78, Rot75, Rot79a, San72, Sch78, Sch79, Sea79, Sei74, Sen75, Sha79, SW73, Sie79, Sin77, Sin79, Siz77, Smi77, Smi79]. matrices [Tam75, Tau72, Tau75, TF71, Tho75a, Tho79b, Tho79c, Var77, Ves79, Wal75, Was77, Wat76, Wil75b, Wim74, Win71a, YI79, Yar78, You79, Zel79a, Zen77, dO74, vK70, vdS79, BVW78, RR75]. Matricial [Bar73b, Deu70, Deu73, DM78, Vit79, Mly77]. Matrix [BM76, Bos75, CF72, FL74a, Fer78a, GLR78a, GLR78b, IM79, Lan71, Lan73, Mey78a, O’R77, Olk79, Tan75, Var75, Vet75, Web74, Aba79, AM76, Arn76, AY73, BK79a, Bal78b, Bar75, Ber75, Ber74, BBI73, BP76, BK78, Bod79, BG74b, BG76a, Bra79, Bre75, But72, Car76, CD79, Cha75, Coo75, Cro78, DD76, Dau78, Djo75, Dow71, Eis73, Els76b, ES73, Eng73, Fer78b, FG77, Fre73, Ger77, Gil77, GR79, GG77, GW76, GZ79b, Hay70, Hed75, Hil77, Hil78, Hoc79, HMW79, HW79c, HM75, Joh77b, Joh79, JB70, Kan78b, Kat72, Kel71, KS78, LM77, Lan77c, Leh71, MN70, MN77, MSH77, MA79, Mey78b, MOV78, Naz77, Ng76, Nwo78, OS72, Pao73a, Pao73b, Par75, Ple77, Por75, PBA73, Reh77, Ros76a, Rot79b, Rys70, Sai72, Sau77, SS76]. matrix [Sch77, Smi71, Smi74, Sta78, Sta75, Tho76a, Tho71, TW72, Tho75b, Tho76c, TW78, TC78, Tur77, Uhl74, Usm71, Wat72, Wer73, Wer79, Wil77, dBG78, dO72, Wer73]. Matrix-theoretic [CF72]. matrix-valued [MOV78]. matrix-vector [TC78]. matrizen [Sch78, Wim78]. matrizenräumen [Bod79]. matroid [BW71, Cun77]. matroids [Gra75, Mau75]. Maximal [Cou73, AS79, Est79, Pal72]. Maximum [PR77, FG77, Mau75, Ptá79]. maximum-rank [Mau75]. mean [Har75a, Mey78b]. means [Pal72]. measure [Chr79]. meeting [Ano77b]. Mendelsohn [Kim79]. Method [Tan75, GJ70, HW79b, LNPU73, Moh76, Mül73, Nat73, Naz75, RBI70]. Methods [HW79c, Van71, Aba79, HMW79, Kos78, Law75, Neu78, Sto79]. Minimal [CV73a, CV73b, Kov77, LS76, Zen77, HT73, PJ75]. minimality [Mly77]. minimization [Naz77, Sto79]. minimum [Ben70, Hed75, Mau75, PR77]. minimum-term-rank [Mau75]. Minkowski [Kan78a]. minors [BW71, Moy77, Pan79a, Tho75b, Tho79b, Hed75]. Mirsky [Lev72b]. mixed [Bal71, Hoc79, Rag78]. mixed- [Bal71]. model [Sie79]. Modified [Ste77, LNPU73]. modularity [Sai72]. modules [Car72]. modulo [BM76]. moment [Par73]. monic [GR79]. monoids [Pra76]. Monotone [Man74, AT76]. monotonic [Fun79]. Monotonicity [HM70, BG72, BP76, Car76, Wer79]. Moore [Cam77, HM75, wK77a, MSH77, Smi77, Smi78]. Motzkin [RS71]. MP [PB79]. multidimensional [AG75]. Multigradients [Hou71a]. Multilinear [Rob78b, Mer78]. multiples [BC77]. multiplication [Hol79, Win71a]. multiplications [Cla78a]. Multiplicative [Fri75, AS79]. Multiplicativity [Joh77b]. multivariate [Sty73].

Nachweis [Met73]. natural [Tho76a]. Near [Qua75]. Nearly [Lan71, Lan73, Min72, BH79, FT79]. Necessary [Far77]. negative [Cha75, Els76b, KS78, Kwo77, PR77]. negativity [Met73]. network [AT76, HP79b, Tch79]. networks [Fla74]. Neumann [Tan75]. Neumann-Type [Tan75]. nichtnegativen [Wer73]. nichtnegativer [Sch78]. Nilpotence [Dix76]. Non [Bye70, HP78, BEK78, Cha75, Els76b, Joh74c, KS78, Kwo77, Met73, PR77, Rob76]. Non-Archimedian [Bye70]. non-commutative [BEK78]. non-negative [Cha75, Els76b, KS78, Kwo77, PR77]. non-negativity [Met73]. non-positive [Joh74c]. Non-singular [HP78]. nondiagonalizable [LNP70]. Nondifferentiable [MF78]. Nondiscrete [Ptá76]. nonhomogeneous [Har77]. nonlinear [Bal71, MR73, Rob76]. Nonnegative [JGK79, Sie79, AY73, BG74b, BG76a, Els76a, FM79, HW79a, LM78, LM79, Mar70b, Ost74, Ple72, Pul71, Rot75, Rot79b, San72, Sch78, Wer73]. nonnormal [KDW74]. nonpositive [BVW78]. Nonsingular [Hea77, Uhl76, Bot78, Lev72b, Uhl74, Ple77]. Nonzero [Mer77]. Norm [GS79, AS73, Pat76, Rob75, RR73, Rob76, Sau77]. norm-numerical [Sau77]. Normal [DR77, Par73, Ruh75, DGS76, Wil75b]. Normality [Joh76b, Laf79]. norme [RR73, Rob76]. normed [AS73, Flo79]. Norms [Mar73, Var73, You79, Bar73b, Bye70, Deu70, Deu73, Fri79, Fun79, Joh77b, Joh79, Par79, Sch70, Sta78, Vit79, Web74]. Note [Bar76, Coh74, Ste79a, dO74, Bra79, Cal73, Cor75, Duk70, Els76a, Gro77, HT73, wK77b, LM77, MR75, Mly77, Neu79, Ros76b, Tam75, Van76, Var77]. notes [Lon71]. Notice [Ano77a]. number [But72, Kim74, Mur72, Rot79b, FL74a]. numbers [GJ70, Mar73]. Numerical [Bal78b, DD75, GTZ75, Hal79, Van77, AYP79, Fox78, FL74b, GS76, GS77, GS79, Joh76b, MA77, MF78, Sau77]. numerically [GM73, dBG78].

oblique [BK79b]. observations [BT73]. Ochoa [Reh77]. off [BVW78]. off-diagonal [BVW78]. One [Aba79, Fei77, Nic75]. One-dimensional [Fei77]. open [Sie79, Var77]. operation [Lon79]. operations [Hil78]. operator [AT76, MR75, PT79b, Spa78]. Operators [Var73, BM79, BW70, Eis73, Fei77, FW74, HFP76, Hoc70, Hol78, MP79, Par79, Van76]. optimal [BD78, Zel79b]. optimization [Pad76]. Optimum [Laf75]. orbital [Gro77]. Orbits [Hof76b]. Order [HM78, BG74a, GK76, HMW79, Kwo77, Mat75, Rob71, SS79, WHG79]. ordered [Els74, Van76]. orderings [Neu78, Pan78]. orders [BK73]. ordinary [FL74b, Nat73, Was77, Lan77a]. orthant [Fun79]. Orthogonal [And74, Kap75, vdSV79, BK79b, Bar75, Wii78]. orthostochastic [AYP79, Hei78]. Oscillation [KP74]. oscillatory [Mar70a]. Ostrowski [OP79]. other [Coh78].

P [BVW78]. packing [Moo74]. Padé [Hou71b]. Pair [Uhl76, And74, Dav70]. Pairs [Uhl79, BW79, Djo78, FO78, Ng76, TC78, Wat76]. Pall [Cor75]. Parallel [Tch79]. parameter [Sta75]. parameters [AH79]. Parametrized [Wer70]. paramount [LS76]. part [Cai73, Joh73]. partial [Neu78]. partially [Els74, Joh74c, Van76]. partially-ordered [Els74]. partitioned [Bar76, De 71, Eis73, HM75, Joh71b, LNP70, LNP71, Mey71, Mih71, vdS79]. Partitioning [Pal72]. partitionings [Hof72]. partitions [Cor75, Por75]. passage [Mey78b]. patterned [Kan78b]. patterns [JLR79]. Pencil [Uhl76, Van79, AY75, Tho76c]. pencils [Oko79, TW72]. Penrose [Cam77, HM75, wK77a, MSH77, Smi77, Smi78]. Perfect [Bar78]. Periodic [Don71]. permanent [Hed75, LO78, Nic75, PJ75]. permanent-like [PJ75]. Permanental [BW79, Bax78, Gib78]. permanents [Joh74b]. permutation [Cru75, Kap75, Ong76]. permutations [FT79, Kap72]. Perron [BS75, RR73, Rob78a]. persymmetric [Zel79b]. Perturbation [Ste79a, Ste79b]. perturbations [Fei77, Sch79, Tho76b]. Perturbed [Pai74, Rob73]. Pfaffians [Kah71]. Piecewise [Eav73]. Pivot [Coh74, Kos78]. plane [BC75]. plusieurs [Rob76]. point [Rob76]. points [MF78]. polyhedra [BG76b, GP79, Pad76]. polyhedral [HFP76]. polyhedron [Cun77]. polynomial [Bar71, Ben70, BCL75, Bra79, Cla78b, Lan77c, Leh71, MB78, Tho76c, Vit79]. Polynomials [GLR78a, GLR78b, Bar70, Bar75, Bax78, Ber74, Deu70, Deu73, Deu78, GR79, HT73, KP74, MB78, RW77]. polytope [BL78, Cru79]. Positive [Cho75b, Djo76, FW74, GV79, Tam76, And79a, Bal70, Cho75a, Cry73, Cry76, De 71, Far77, Far76, GTZ75, Hir74, HW79b, Joh73, Joh74c, JLR79, KP74, Kel71, Mee76, Moy77, Nic79, Pan79a, Pan79b, Tam75, Wer73]. positiven [Wer73]. positiver [Hir74]. positives [Far76]. positivity [MS78, NP79]. potent [Kim75]. pour [Mor76, RR73, Rob76]. Power [Joh79, AY73]. Powers [Hua78, DD76, Kwo77, MR76, MA77, You79]. Preface [Ste79c]. prescribed [Bax78, BC73]. preserve [Bot78, Hil73, LO78, MW79, Wat76]. preserved [Kwo77]. preserving [BT73, Hor78a]. primal [CGG77]. Principal [Tho72, Tho79b, BW71, McD78, Moy77, Pan79a, Tho76c]. probability [Har75a]. Problem [Ste79b, CH75, Dat78b, Dav76, DR76, FW74, GR79, Kan78b, Kos78, Lev72b, Lyl78, MN77, Mor76, Mur72, Pan78, Ptá79, Reb74, Roi78, Wil77]. problème [Mor76]. Problems [Fri75, Fri77, CGG77, Els79, Hal76a, Hoc75, Kan78a, Mer75, Moh76, Nat73, Pad76, Tur77, Usm71]. procedure [Ban79]. processes [Har77]. product [O’R77, Pat76, Tho76a, Tho71, Win71b]. production [Har77]. Products [Bal70, Bal75, Bal78c, Cat78, DM79, Ell79, Gib76, GHR76, HK76, And79a, Ara79, Ber75, Dav70, DGS76, ES73, Heu73, Kat75, eK73, Oko79, Par75, Sin79, Sty73, Hir74]. Produkte [Hir74]. program [Rob73]. programming [Bal71, CGG77, DR76, KR73, Man79, Par70, RBI70]. Projection [vdSV79, GJ70]. projectors [BK79b]. proof [Cru75, Dat78a, Fla74, Ost71]. proofs [HP79b, Ptá76, vLvEB78]. proper [Law75]. Properties [CM75, Usm71, Ara79, Cru77, Cry76, GS79, Hay70, Hei78, HM78, How78, KP74, eK73, Mar79b, MSH77, Mur72, NS79, Zel79a]. property [BK73, RS71, Tho70]. propres [Mor76]. proving [AG75, Kah71]. pseudo [CF72, JO75, Paz75]. pseudo-convexity [CF72]. pseudo-isolated [JO75]. pseudo-Markov [Paz75]. pseudoconvex [Fer78a]. Pseudoinverse [CM75, Gre73, PBA73, Tho76a]. pseudoinverses [Deu71]. Putnam [DGS76].

Quadratic [Put79, Tur79, Uhl79, CF72, Dau78, Dav79, Far77, Par70, Sto79]. quasi [CF72, Joh77a]. quasi-convexity [CF72]. quasi-elementary [Joh77a]. quasiaccretive [MN79]. quasidiagonalization [Laf77]. Quasimonotonicity [Els74]. Quasimonotonie [Els74]. quasiorder [Bod79]. Quasiordnung [Bod79]. quaternion [DB71, Smi74]. question [Var77]. quotient [Ost71].

R. [Fla74, Min79]. radii [GS79]. radius [Els76b, Fer78b, GTZ75, MA77, Mül73]. random [SD77]. Rang [Far76]. range [Bal78b, Fox78, Joh76b, MF78, Sau77]. ranges [AYP79, CL74, GS76, GS77, Loe76]. Rank [vdSV79, Bea70, CH77, CF79, How78, Kru77, Mau75, Rot79b, Sto79, Tho76b, Wes72, Far76]. rank- [Sto79]. Ranked [PM79]. ranks [Bra70]. rate [BM79]. rational [AY73, CP79, Dow71, Lam75, Tau75]. rationaler [Wim78]. Räumen [Els74]. Rayleigh [Van71]. Real [Gib78, Maj79, Uhl73, Uhl76, AY75, BC77, Bos75, Cai73, DR77, Hir74, Loe76, Neu79, Ng76, Pan79b, Smi74, Tho71, Uhl74, Ves79]. realisierungen [Wim78]. Rearrangement [Mar75]. reciprocal [LNP71]. reconstruction [dBG78]. rectangular [Neu76, Neu78]. Recurrence [Par76]. Recurring [Uhl79, Var76b]. recursions [Mat75]. recursively [Ara79]. reducible [BH79]. Reduction [Hay70, Kan78b]. reellen [Hir74]. Referee [Ano76g, Ano77i, Ano78k, Ano79l]. referees [Ano73e]. reflections [DM79]. reflexion [Eav73]. Reflexive [DF75, Tan75, Deu71, HK76]. Regular [Eng73]. Reich [OP79]. related [AT76, BG74a, Dav79, Hoc75, Jen73, MR73, Pad76, Tho75a]. Relations [Par75, BK79b, CP79, GS76, GS77, TW72, Wil70]. relationship [Sin79]. Relaxation [Don71, Mül73]. relaxationsverfahren [Mül73]. remark [Cha70, Pao73a]. Remarks [JB70, Kat72, Mey70]. removing [Cru79]. Rényi [KR79b]. representation [AS73, Gil77, Mih71]. Representations [Mey71, IM79, Ros76a]. Representing [Dav70]. Resolvent [GLR78b]. respect [DR77]. Restoring [vdSV79]. restricted [Tho76b]. Results [Tam76, BG74a, Bre74, Har75a, Kat72, MN70, Mar75, SS79, Tho75a]. retractions [Eav73]. retrocirculant [Smi78]. reverse [BG74a, SS79, WHG79]. reverse-order [SS79]. Riccati [Mey78a]. Riemann [Gil77]. ring [DM79]. rings [BG78, Chi77, Est79, Gus79, LM77, McD78]. Robert [Bod79]. Robert’schen [Bod79]. role [Tau72]. root [BG74b, BG76a, Dat78b]. root-separation [Dat78b]. Roots [Hea70, Gib78, HW79b, JGK79, MN79, Ost74, TW72, dO74]. Rosenberg [Rob78a]. Roth [Gus79]. Routh [Dat78b]. row [BC73, Har78, HP78, Joh74a, Kap72]. row-space [Har78]. rows [Mar79d].

S [Ano77a, Ste79c, Wil75a, BVW78]. S. [Ano78a]. satisfied [Far77]. satisfying [Lam75]. Scalar [BCL75, Heu73]. scaling [Ban79]. scheme [Ban79]. schemes [AG75, Tur77]. Schur [And79b, BJ78, Cot74, Dat78b, Hay70, MM72b, Mer75, Ost71, Wat72]. Schwarz [Mar79a]. scientific [Gal78]. Scott [Min79]. second [Mat75]. Self [BF76, BIL78, Gra75]. Self-dual [BF76, BIL78, Gra75]. Semi [Deu71, AY75, Djo76, Far76, Smi76]. semi-définies [Far76]. semi-definite [Djo76]. semi-definiteness [AY75]. Semi-inverses [Deu71]. semi-linear [Smi76]. semicircle [Arn76]. semiconvergent [Ple76]. semidefinite [Cai73, Cho75b, De 71, Far76]. semigroup [Mar70b]. semigroups [But72, HK76, Put79, Siz77, Web74]. semilinear [Djo78]. seminorms [Fer78b]. semipositive [BW78]. semisimple [Ion76, LC70]. Separable [Rob71]. separation [Dat78b]. sequence [NS79]. sequences [Uhr77]. series [Sta75]. serve [Heu73]. sesquilinear [Fox78]. sesquillinear [Dav76]. set [BD78, Far77, Rob73, Zel79b]. sets [Eck75, Eng73, Sha79, Wil70]. several [Rob76]. Sharpness [ASV79]. Shifted [DT71]. Shorted [MP79]. Sign [JLR79, Gre74]. sign-symmetric [Gre74]. similar [Kel71, Par75]. Similarity [McD78, BW70, Djo73, DR77, Hil78]. similitudes [Tho75c]. simple [Rad75]. simplex [GM73, Moh76]. simplification [Was77]. Simultaneous [Laf77, MN77, Sha79, Uhl73]. Singular [Lan71, Lan73, MS78, Ruh75, Ste79a, Tho75b, Tho79c, Van79, Var75, Bot78, Cam79, HP78, wK77a, wK77b, Nat73, Smi79, Tho70, Tho71, Tho72, Tho76b]. Size [Coh74]. skew [And74, AHS76, Djo75, Siz77, Tho79b]. skew-Hermitian [Djo75]. skew-symmetric [And74, AHS76]. skewfields [Ell79]. Smallest [Var75]. solution [Hal79, HMW79, LNPU73, Man79, Rob73]. Solutions [Hea77, Kov77, Vet75, Bar77, Cam79, Dau78, FL74b, FM78, Kov76, Mat75, Mey78a, Mur72, PM79, WHG79]. Solvability [Por76, BBI73]. Solving [Spa78, HW79c, Kos78]. Some [Ara79, AY73, BT73, Bre74, CP79, Cru77, Cry76, Fun79, eK73, Lon71, MN70, Mey70, Neu78, NS79, Por75, Tam76, vLvEB78, Ach78, Bal78a, Bal78b, Els79, HM78, LNP71, Zel79a, vK70]. SOR [OP79]. Souriau [Gre73]. space [Har78, Heu73, Sto79, Vid76]. Spaces [Wes72, AS73, BF76, Bod79, Bra70, Els74, FO78, Flo79, Kov74, Kov75, Moo74, Put79, Qua75, Van76]. spanning [Mur72]. special [Joh74a, Reh77, vK70]. specified [KS78]. spectra [BT73, BM79, Doo70, You79]. Spectral [GLR78a, GLR78b, Har78, BEK78, CH75, Doo75, DC79, Els76b, Fer78b, GR79, GW76, Hoc75, Joh79, Mül73, dBG78, FL74a]. spectral-radius [Mül73]. Spectrum [Kov74, Kov75]. spektralradius [Mül73]. spheres [Tur79]. spherical [KW79, Mer78]. spline [FM78, LS73, Nat73, SS73]. splines [Lee75, SS73]. splitting [Neu76, Man74]. splittings [Law75, NP79, Ple76]. spread [Deu78]. square [MN79, Pao73a, Pao73b, Kap72]. squares [Cia70, Djo76, LL78, LNPU73]. Stability [Dat78c, AG75, Cro78, Dat78a, GK76, Joh76a]. Stabilization [Yar78]. stable [BW78, Cai73, GM73, HW79b, Ros76a, dBG78]. standard [RR73]. star [MR75]. state [Sta75]. statistical [Sty73]. statistics [Kru77]. Stein [Rob78a, Wim75]. Stieltjes [MW79]. stochastic [Ach78, BG75, BW79, BC75, BG76b, Chr79, Fie72, Gib76, Gib78, Hed75, Joh71a, Joh74a, Joh77a, Lev72b, MT79, Sen75, Sin77, Sin79, Wal75]. strict [Neu79]. strictly [Fan72, MN79]. Structure [And78, Tho75c, BL75, GR79, Vid76]. Structures [Vet75]. study [Tau72]. subadditive [Zen77]. subalgebras [Cou73]. subcones [Tam78]. subdominant [Ost74, Sch78]. subdominanten [Sch78]. subfields [LM77]. submatrices [CH77, LNP70, Tho72]. suboptimization [RBI70]. subpencil [Tho76c]. Subproper [Neu76]. subsemigroups [KR79a]. subspace [FHL77, Gro77]. subspaces [NS79, Pal72]. sufficient [Far77]. sum [Afr74]. sum-symmetric [Afr74]. sums [Ach78, BC73, Cia70, Djo76, HP78, Joh76a, LL78, Oko79, Par79, PR77, Sin77, TF71, Tho75b]. surfaces [Gil77]. Survey [Uhl79]. Symmetric [Maj79, Tur77, Uhl73, Uhl76, Afr74, And74, AHS76, Ben70, Ber74, CB76, DB71, Dia79, FT79, Fle76, Gre74, Hor75, JGK79, Joh73, Kap72, LO78, MM72a, Ng76, Pat76, RW77, Smi79, Tau72, Tho79b, Tho79c, Uhl74, Zel79b, dOdS79]. Symmetrien [Bod79]. Symmetries [Bod79]. symmetrized [Mer77]. symmetry [BW70, Cha79]. System [Van77, LNPU73]. Systems [FL74b, GV79, Fan75, Neu78].

Table [Hou71b]. taking [Kwo77]. Taussky [RS71]. Taylor [Sta75]. tensor [Dav70, How78, eK73, Tho76a]. tensors [Dia79, Mer77, dOdS79]. term [Mau75]. terms [FMKL79]. Ternary [Coh78]. th [HW79b, JGK79, Kap72]. their [Doo70, FMKL79, Sta78]. Theorem [ASV79, Lan77a, Lan77b, Uhl79, Bal71, BEK78, Bar79, Ber75, Ber76, BBI73, Dav77, DGS76, Fan75, Fer78b, Fu75, GK76, HP79b, HP79a, Jen73, Joh76a, Kah71, Kan78b, KR73, Kim79, Lev72a, Loe71, Mau75, OP79, Pow76, Sil70]. Théorèmes [Rob78a]. Theorems [Eck75, Kat75, Bri78, Dat78b, Gus79, Har75a, Hof76a, Joh71b, Rob78a, Ros76a, Var76b, Wim74, Wim75]. theoretic [CF72, KR79a]. théorie [RR73]. theory [AT76, BS75, BBI73, Bre74, CH76, DD75, FW74, Har75a, Hoc75, Joh71a, Kat72, Lyl78, Pul71, RR73, Sha77, TW72, Wii78, vLvEB78]. Three [Cro78, Kru77, AH79, BC75, BK73, Cou73, Fox78]. three-by-three [BK73]. Three-way [Kru77]. timetable [SS76]. Title [Ano70c, Ano71c, Ano73f, Ano73g, Ano79m]. Toeplitz [AHS76, CPW74, FMKL79, GT79, LM79]. Topological [Hei78]. Total [GP79, Hof76a, Reb74, Met73]. totally [Cry73, Cry76, KP74, LM79, Mar70b]. Totalnichtnegativität [Met73]. tournament [dO74]. Trace [PT79a]. Transcendental [Hou71a]. transform [Fri79]. transformation [Deu71, Djo73, FHL77, Mer71, Rob78b, Smi76]. Transformations [Gen75, Bea70, CL74, Cla78a, Dav79, DF75, Duk70, FO78, Hil73, Kim74, LO78, Loe76, OS72, Ong76, Wes72]. transforms [Fan72]. transition [Har75a, Sta75]. transportation [Reb74]. Transposing [Hil78]. Treatment [Van77, BH79]. Triangular [Par76]. triangularization [Sha79]. Triangularizing [Siz77]. tridiagonal [Bar79, RS74, Ves79]. trilinear [Kru77]. triviality [Cha79]. Truncated [Sta75, BL78]. Two [BK79b, Fan75, Har75a, Mar79b, Mer75, Sto79, Uhl73, AY75, Bar70, Ell79, Ful73b, Ful79, Kat72, Kwo77, Wii78]. Type [Tan75, Cha73, Olk79]. types [BP76, Cro78, Heu73].

unbounded [Cun77, GK76]. unconstrained [Naz77]. Ungleichungen [Els74]. Unified [Naz77, BH79]. uniformly [NS79]. unimodular [And78]. unimodularity [Hof76a, Reb74, TC78]. Uniqueness [Man79, Kru77]. unitarily [Web74]. Unitary [BW70, Hol78, IM79, MG70, MN70, TW78]. univalence [GZ79a]. Unsymmetric [GV79]. upon [vK70]. using [Hay70, Hil77]. utilizing [MSH77].

V [SS73]. valeurs [Mor76]. Value [Var75, Nat73, Tho75b, Usm71]. valued [MSH77, MOV78]. Values [Ruh75, Ste79a, MA79, Tho70, Tho71, Tho72, Tho76b, Tho79c]. Vandermonde [Van77]. variables [Rob76]. Variation [vK70]. Variations [Mar79a, Sil70]. Vector [Moo74, Vet75, Bra70, Bre74, FO78, Flo79, Qua75, RR73, Rob76, TC78, Van76]. vectorielle [RR73, Rob76]. vectors [FT79, O’B75, dO72]. verallgemeinerter [Mül73]. version [Arn76]. vertex [Cru79]. via [Naz77]. vibrations [Wim74]. VII [Ano77b]. Volume [Ano77d, Ano75a, Ano75b, Ano75c, Ano76a, Ano76b, Ano76c, Ano77c, Ano77e, Ano78c, Ano78d, Ano78e, Ano78f, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e]. volumes [Ano78b, Ano79f, Ano79m].

Wachspress [AH79]. Waerden [Lon71]. way [Kru77]. Weak [CM78, RR73]. weakly [Gre74]. weighted [PR77]. whenever [Far77]. where [LV75, LV76]. which [Hed75, Heu73, Hil73, Kap72, Kwo77, MW79, Sin77]. whose [Hof77, Joh73]. Widths [Pin79]. Wigner [Arn76]. Word [Paz75]. work [Gal78]. works [Sch77].

Z [Coh78]. Zeros [Hou71a, Bar71, Cla78b, Deu70, Deu73, Rya79, Vit79]. zilch [Hea75, Hea73].

References

Abaffy:1979:OCD

[Aba79] József Abaffy. One class of dual matrix methods. Linear Algebra and its Applications, 27(??):93–102, October 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

         sciencedirect.com/science/article/pii/002437957990034X.

Achilles:1978:DSM

[Ach78] Eva Achilles. Doubly stochastic matrices with some equal diagonal sums. Linear Algebra and its Applications, 22(??):293–296, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

         tronic). URL http://www.sciencedirect.com/science/article/
         pii/0024379578900794.

Afriat:1974:SSM

[Afr74] S. N. Afriat. On sum-symmetric matrices. Linear Algebra and its Applications, 8(2):129–140, April 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379574900512.

Abarbanel:1975:PSM

[AG75] S. Abarbanel and D. Gottlieb. On proving stability of multidimensional difference schemes. Linear Algebra and its Applications, 11(3):247–250, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379575900245.

Avdelas:1979:JWP

[AH79] G. Avdelas and A. Hadjidimos. Jordan–Wachspress parameters in three dimensions. Linear Algebra and its Applications, 24(??):251– 261, April 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379579901629.

Andres:1976:DCS

[AHS76] T. H. Andres, W. D. Hoskins, and R. G. Stanton. The determinant of a class of skew-symmetric Toeplitz matrices. Linear Algebra and its Applications, 14(2):179–186, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379576900227.

Anderson:1976:MKL

[AM76] Brian D. O. Anderson and John B. Moore. A matrix Kronecker lemma. Linear Algebra and its Applications, 15(3):227–234, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900288.

Andrew:1973:ECM

[And73] Alan L. Andrew. Eigenvectors of certain matrices. Linear Algebra and its Applications, 7(2):151–162, April 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379573900499.

Anderson:1974:ODD

[And74] B. D. O. Anderson. Orthogonal decomposition defined by a pair of skew-symmetric forms. Linear Algebra and its Applications, 8(1):91– 93, February 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379574900123.

Anderson:1978:SCU

[And78] Brian D. O. Anderson. Structure of a class of unimodular matrices. Linear Algebra and its Applications, 22(??):167–174, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957890068X.

Ando:1979:CCM

[And79a] T. Ando. Concavity of certain maps on positive definite matrices and applications to Hadamard products. Linear Algebra and its Applications, 26(??):203–241, August 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579901794.

Ando:1979:GSC

[And79b] T. Ando. Generalized Schur complements. Linear Algebra and its Applications, 27(??):173–186, October 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579900405.

Anonymous:1970:AI

[Ano70a] Anonymous. Author index. Linear Algebra and its Applications, 3 (4):491–492, October 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/002437957090039X.

Anonymous:1970:EB

[Ano70b] Anonymous. Editorial Board. Linear Algebra and its Applications, 3(1):ifc, January 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379570900212.

Anonymous:1970:TI

[Ano70c] Anonymous. Title index. Linear Algebra and its Applications, 3 (4):493–494, October 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379570900406.

Anonymous:1971:AI

[Ano71a] Anonymous. Author index. Linear Algebra and its Applications, 4(4): 421, October 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379571900139.

Anonymous:1971:EB

[Ano71b] Anonymous. Editorial Board. Linear Algebra and its Applications, 4(1):ifc, January 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379571900231.

Anonymous:1971:TI

[Ano71c] Anonymous. Title index. Linear Algebra and its Applications, 4(4): 424, October 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379571900140.

Anonymous:1972:EB

[Ano72] Anonymous. Editorial board. Linear Algebra and its Applications, 5(1):ifc, January 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379572900122.

Anonymous:1973:AIa

[Ano73a] Anonymous. Author index. Linear Algebra and its Applications, 6(??):337, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379573900311.

Anonymous:1973:AIb

[Ano73b] Anonymous. Author index. Linear Algebra and its Applications, 7(4): 367, October 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0024379573800096.

Anonymous:1973:EBa

[Ano73c] Anonymous. Editorial Board. Linear Algebra and its Applications, 6(??):ifc, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379573900013.

Anonymous:1973:EBb

[Ano73d] Anonymous. Editorial board. Linear Algebra and its Applications, 7(1):ifc, January 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379573900335.

Anonymous:1973:LR

[Ano73e] Anonymous. List of referees. Linear Algebra and its Applications, 7(4):366, October 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0024379573800084.

Anonymous:1973:TIa

[Ano73f] Anonymous. Title index. Linear Algebra and its Applications, 6 (??):339–340, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379573900323.

Anonymous:1973:TIb

[Ano73g] Anonymous. Title index. Linear Algebra and its Applications, 7 (4):369–370, October 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0024379573800102.

Anonymous:1974:AJ

[Ano74a] Anonymous. Addendum (July, 1973). Linear Algebra and its Applications, 8(2):104, April 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379574900470.

Anonymous:1974:EB

[Ano74b] Anonymous. Editorial board. Linear Algebra and its Applications, 8(1):ifc, February 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379574900019.

Anonymous:1975:AIVa

[Ano75a] Anonymous. Author index to volume 10. Linear Algebra and its Applications, 10(3):291–292, June 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379575900762.

Anonymous:1975:AIVb

[Ano75b] Anonymous. Author index to volume 11. Linear Algebra and its Applications, 11(3):291, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379575900282.

Anonymous:1975:AIVc

[Ano75c] Anonymous. Author index to volume 12. Linear Algebra and its Applications, 12(3):293, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379575900518.

Anonymous:1975:EBa

[Ano75d] Anonymous. Editorial board. Linear Algebra and its Applications, 10(1):ifc, February 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379575900889.

Anonymous:1975:EBb

[Ano75e] Anonymous. Editorial Board. Linear Algebra and its Applications, 11(1):CO2, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/002437957590110X.

Anonymous:1975:EBc

[Ano75f] Anonymous. Editorial Board. Linear Algebra and its Applications, 12 (1):ifc, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379575901202.

Anonymous:1976:AIVa

[Ano76a] Anonymous. Author index to volume 13. Linear Algebra and its Applications, 13(3):291–292, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379576901038.

Anonymous:1976:AIVb

[Ano76b] Anonymous. Author index to volume 14. Linear Algebra and its Applications, 14(3):293, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379576900744.

Anonymous:1976:AIVc

[Ano76c] Anonymous. Author index to volume 15. Linear Algebra and its Applications, 15(3):294, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379576900355.

Anonymous:1976:EBa

[Ano76d] Anonymous. Editorial board. Linear Algebra and its Applications, 13(1–2):ifc, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379576900367.

Anonymous:1976:EBb

[Ano76e] Anonymous. Editorial Board. Linear Algebra and its Applications, 14 (1):ifc, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379576900562.

Anonymous:1976:EBc

[Ano76f] Anonymous. Editorial Board. Linear Algebra and its Applications, 15 (1):ifc, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379576900756.

Anonymous:1976:RL

[Ano76g] Anonymous. Referee list. Linear Algebra and its Applications, 15(3): 293, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379576900343.

Anonymous:1977:AHA

[Ano77a] Anonymous. Alston S. Householder Award (1977) — final notice. Linear Algebra and its Applications, 17(3):189, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957790060X.

Anonymous:1977:AGV

[Ano77b] Anonymous. Announcement of the Gatlinburg VII meeting. Linear Algebra and its Applications, 16(1):93, ???? 1977. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379577900234.

Anonymous:1977:AIVa

[Ano77c] Anonymous. Author index to volume 16. Linear Algebra and its Applications, 16(3):293, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379577900118.

Anonymous:1977:AIVb

[Ano77d] Anonymous. Author index to volume 17. Linear Algebra and its Applications, 17(3):293, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379577900684.

Anonymous:1977:AIVc

[Ano77e] Anonymous. Author index to volume 18. Linear Algebra and its Applications, 18(3):293, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379577900593.

Anonymous:1977:EBa

[Ano77f] Anonymous. Editorial board. Linear Algebra and its Applications, 16 (1):ifc, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/002437957790012X.

Anonymous:1977:EBb

[Ano77g] Anonymous. Editorial board. Linear Algebra and its Applications, 17 (1):ifc, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379577900374.

Anonymous:1977:EBc

[Ano77h] Anonymous. Editorial Board. Linear Algebra and its Applications, 18 (1):ifc, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/002437957790074X.

Anonymous:1977:RLA

[Ano77i] Anonymous. Referee list: August 1977. Linear Algebra and its Applications, 18(3):292, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379577900581.

Anonymous:1978:AHA

[Ano78a] Anonymous. The Alston S. Householder award. Linear Algebra and its Applications, 21(1):1, July 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379587901959.

Anonymous:1978:AIVc

[Ano78b] Anonymous. Author index for volumes 1–20. Linear Algebra and its Applications, 20(3):275–292, June 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379578900241.

Anonymous:1978:AIVa

[Ano78c] Anonymous. Author index to volume 19. Linear Algebra and its Applications, 19(3):293, ???? 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379578900162.

Anonymous:1978:AIVb

[Ano78d] Anonymous. Author index to volume 20. Linear Algebra and its Applications, 20(3):271, June 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/002437957890023X.

Anonymous:1978:AIVd

[Ano78e] Anonymous. Author index to volume 21. Linear Algebra and its Applications, 21(3):293, September 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379578900903.

Anonymous:1978:AIVe

[Ano78f] Anonymous. Author index to volume 22. Linear Algebra and its Applications, 22(??):297, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379578900800.

Anonymous:1978:EBa

[Ano78g] Anonymous. Editorial board. Linear Algebra and its Applications, 19 (1):ifc, ???? 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379578900010.

Anonymous:1978:EBb

[Ano78h] Anonymous. Editorial Board. Linear Algebra and its Applications, 20(1):ifc, March 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379578900253.

Anonymous:1978:EBc

[Ano78i] Anonymous. Editorial Board. Linear Algebra and its Applications, 21 (1):ifc, July 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379587901947.

Anonymous:1978:EBd

[Ano78j] Anonymous. Editorial Board. Linear Algebra and its Applications, 22(??):ii, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379578900502.

Anonymous:1978:RLJ

[Ano78k] Anonymous. Referee list: June 1978. Linear Algebra and its Applications, 22(??):298, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379578900812.

Anonymous:1979:AIVa

[Ano79a] Anonymous. Author index to volume 23. Linear Algebra and its Applications, 23(??):281–282, February 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579901101.

Anonymous:1979:AIVb

[Ano79b] Anonymous. Author index to volume 24. Linear Algebra and its Applications, 24(??):289, April 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579901654.

Anonymous:1979:AIVc

[Ano79c] Anonymous. Author index to volume 25. Linear Algebra and its Applications, 25(??):299, June 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579900259.

Anonymous:1979:AIVd

[Ano79d] Anonymous. Author index to volume 26. Linear Algebra and its Applications, 26(??):297, August 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579901848.

Anonymous:1979:AIVe

[Ano79e] Anonymous. Author index to volume 27. Linear Algebra and its Applications, 27(??):279, October 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579900478.

Anonymous:1979:AIVf

[Ano79f] Anonymous. Author index to volumes 23–28, 1979. Linear Algebra and its Applications, 28(??):305–309, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901411.

Anonymous:1979:EBa

[Ano79g] Anonymous. Editorial Board. Linear Algebra and its Applications, 23(??):ii, February 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379579900879.

Anonymous:1979:EBb

[Ano79h] Anonymous. Editorial Board. Linear Algebra and its Applications, 25 (??):ii, June 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379579900016.

Anonymous:1979:EBc

[Ano79i] Anonymous. Editorial Board. Linear Algebra and its Applications, 26(??):ii, August 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379579901666.

Anonymous:1979:EBd

[Ano79j] Anonymous. Editorial Board. Linear Algebra and its Applications, 28(??):ii, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379579901113.

Anonymous:1979:LLA

[Ano79k] Anonymous. Letters in linear algebra. Linear Algebra and its Applications, 26(??):1, August 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379579901678.

Anonymous:1979:RL

[Ano79l] Anonymous. Referee list. Linear Algebra and its Applications, 28 (??):313, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379579901435.

Anonymous:1979:TIV

[Ano79m] Anonymous. Title index to volumes 23–28, 1979. Linear Algebra and its Applications, 28(??):310–312, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579901423.

Arazi:1979:SPH

[Ara79] Benjamin Arazi. Some properties of Hadamard matrices generated recursively by Kronecker products. Linear Algebra and its Applications, 25(??):27–39, June 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/002437957990003X.

Arnold:1976:DVW

[Arn76] Ludwig Arnold. Deterministic version of Wigner’s semicircle law for the distribution of matrix eigenvalues. Linear Algebra and its Applications, 13(3):185–199, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379576900951.

Arms:1973:RNF

[AS73] R. J. Arms and O. Shisha. A representation of the norm in finitedimensional normed linear spaces. Linear Algebra and its Applications, 7(2):173–174, April 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379573900517.

Atkinson:1979:MMC

[AS79] M. D. Atkinson and N. M. Stephens. On the maximal multiplicative complexity of a family of bilinear forms. Linear Algebra and its Applications, 27(??):1–8, October 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579900260.

Anderson:1979:EKT

[ASV79] N. Anderson, E. B. Saff, and R. S. Varga. On the Eneström–Kakeya theorem and its sharpness. Linear Algebra and its Applications, 28 (??):5–16, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379579901137.

Anderson:1976:CMO

[AT76] William N. Anderson, Jr. and George E. Trapp. A class of monotone operator functions related to electrical network theory. Linear Algebra and its Applications, 15(1):53–67, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379576900781.

Au-Yeung:1973:SIR

[AY73] Yik-Hoi Au-Yeung. Some inequalities for the rational power of a nonnegative definite matrix. Linear Algebra and its Applications, 7(4):347–350, October 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0024379573800059.

Au-Yeung:1975:SDR

[AY75] Yik-Hoi Au-Yeung. On the semi-definiteness of the real pencil of two Hermitian matrices. Linear Algebra and its Applications, 10(1):71–76, February 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379575900968.

Au-Yeung:1979:OMC

[AYP79] Yik-Hoi Hoi Au-Yeung and Yiu-Tung Tung Poon. 3 × 3 orthostochastic matrices and the convexity of generalized numerical ranges. Linear Algebra and its Applications, 27(??):69–79, October 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579900326.

Ballantine:1970:PPD

[Bal70] Charles S. Ballantine. Products of positive definite matrices. IV. Linear Algebra and its Applications, 3(1):79–114, January 1970.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379570900303.

Balas:1971:DTA

[Bal71] Egon Balas. A duality theorem and an algorithm for (mixed-) integer nonlinear programming. Linear Algebra and its Applications, 4(4):341–352, October 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/002437957190005X.

Ballantine:1975:PEM

[Bal75] C. S. Ballantine. Products of EP matrices. Linear Algebra and its Applications, 12(3):257–267, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379575900488.

Balasubramanian:1978:ESE

[Bal78a] K. Balasubramanian. On equality of some elements in matrices. Linear Algebra and its Applications, 22(??):135–138, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900642.

Ballantine:1978:NRM

[Bal78b] C. S. Ballantine. Numerical range of a matrix: some effective criteria. Linear Algebra and its Applications, 19(2):117–188, ???? 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900101.

Ballantine:1978:PIM

[Bal78c] C. S. Ballantine. Products of idempotent matrices. Linear Algebra and its Applications, 19(1):81–86, ???? 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437957890006X.

Bank:1979:ASP

[Ban79] Randolph E. Bank. An automatic scaling procedure for a D’Yakanov– Gunn iteration scheme. Linear Algebra and its Applications, 28 (??):17–33, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379579901149.

Barnett:1970:GCD

[Bar70] S. Barnett. Greatest common divisor of two polynomials. Linear Algebra and its Applications, 3(1):7–9, January 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379570900236.

Barnett:1971:LZC

[Bar71] S. Barnett. Location of zeros of a complex polynomial. Linear Algebra and its Applications, 4(1):71–76, January 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379571900280.

Barker:1973:LFF

[Bar73a] George Phillip Barker. The lattice of faces of a finite dimensional cone. Linear Algebra and its Applications, 7(1):71–82, January 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379573900384.

Barker:1973:MNC

[Bar73b] George Phillip Barker. Matricial norms over cones. Linear Algebra and its Applications, 6(??):175–182, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379573900177.

Barnett:1975:CMA

[Bar75] Stephen Barnett. A companion matrix analogue for orthogonal polynomials. Linear Algebra and its Applications, 12(3):197–208, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900415.

Barnett:1976:NIC

[Bar76] Stephen Barnett. Note on inversion of certain block-partitioned matrices. Linear Algebra and its Applications, 13(1–2):133–137, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900513.

Barker:1977:CSL

[Bar77] George Phillip Barker. Common solutions to the Lyapunov equations. Linear Algebra and its Applications, 16(3):233–235, ????

1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900064.

Barker:1978:PC

[Bar78] George Phillip Barker. Perfect cones. Linear Algebra and its Applications, 22(??):211–221, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379578900721.

Barrett:1979:TIT

[Bar79] Wayne W. Barrett. A theorem on inverses of tridiagonal matrices. Linear Algebra and its Applications, 27(??):211–217, October 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579900430.

Batigne:1978:IGI

[Bat78] Daniel Batigne. Integral generalized inverses of integral matrices. Linear Algebra and its Applications, 22(??):125–134, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900630.

Baxter:1978:EMP

[Bax78] David K. Baxter. The existence of matrices with prescribed characteristic and permanental polynomials. Linear Algebra and its Applications, 22(??):175–189, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379578900691.

Berman:1973:LEC

[BBI73] Abraham Berman and Adi Ben-Israel. Linear equations over cones with interior: a solvability theorem with applications to matrix theory. Linear Algebra and its Applications, 7(2):139–149, April 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379573900487.

Brawley:1973:EMP

[BC73] J. V. Brawley and L. Carlitz. Enumeration of matrices with prescribed row and column sums. Linear Algebra and its Applications, 6(??):165– 174, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379573900165.

Brualdi:1975:EPS

[BC75] R. A. Brualdi and J. Csima. Extremal plane stochastic matrices of dimension three. Linear Algebra and its Applications, 11(2):105– 133, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379575900531.

Bahl:1977:IDM

[BC77] Christina A. Bahl and Bryan E. Cain. The inertia of diagonal multiples of 3 × 3 real matrices. Linear Algebra and its Applications, 18(3):267–280, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379577900568.

Brawley:1975:SPF

[BCL75] J. V. Brawley, L. Carlitz, and Jack Levine. Scalar polynomial functions on the n × n matrices over a finite field. Linear Algebra and its Applications, 10(3):199–217, June 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379575900695.

Brockett:1978:OES

[BD78] Roger W. Brockett and David Dobkin. On the optimal evaluation of a set of bilinear forms. Linear Algebra and its Applications, 19(3):207– 235, ???? 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379578900125.

Beasley:1970:LTM

[Bea70] Leroy B. Beasley. Linear transformations on matrices: the invariance of rank k matrices. Linear Algebra and its Applications, 3(4):407–427, October 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379570900339.

Barker:1978:NCS

[BEK78] G. P. Barker, L. Q. Eifler, and T. P. Kezlan. A non-commutative spectral theorem. Linear Algebra and its Applications, 20(2):95–100, ????

1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900435.

Bender:1970:DSM

[Ben70] Edward A. Bender. The dimensions of symmetric matrices with a given minimum polynomial. Linear Algebra and its Applications, 3(1):115–123, January 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379570900315.

Berkowitz:1974:EES

[Ber74] S. Berkowitz. Extrema of elementary symmetric polynomials of the eigenvalues of the matrix P^{∗}KP + L. Linear Algebra and its Applications, 8(3):273–280, June 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379574900731.

Berkey:1975:GTL

[Ber75] Dennis D. Berkey. A Gerschgorin theorem for linear difference equations and eigenvalues of matrix products. Linear Algebra and its Applications, 11(1):27–40, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379575901159.

Berkey:1976:GTD

[Ber76] Dennis D. Berkey. A Gerschgorin theorem for difference equations — II. Linear Algebra and its Applications, 14(1):87–94, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900653.

Barker:1976:SDC

[BF76] George Philip Barker and James Foran. Self-dual cones in Euclidean spaces. Linear Algebra and its Applications, 13(1–2):147–155, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900537.

Berman:1972:GIM

[BG72] Abraham Berman and Prabha Gaiha. A generalization of irreducible monotonicity. Linear Algebra and its Applications, 5(1):29–38, January 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/002437957290016X.

Barwick:1974:GRO

[BG74a] David T. Barwick and Jimmie D. Gilbert. Generalizations of the reverse order law with related results. Linear Algebra and its Applications, 8(4):345–349, August 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379574900615.

Brauer:1974:BGC

[BG74b] Alfred Brauer and Ivey C. Gentry. Bounds for the greatest characteristic root of an irreducible nonnegative matrix. Linear Algebra and its Applications, 8(2):105–107, April 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379574900482.

Beasley:1975:GDS

[BG75] E. T. Beasley, Jr. and P. M. Gibson. A generalization of doubly stochastic matrices over a field. Linear Algebra and its Applications, 10(3):241–257, June 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379575900725.

Brauer:1976:BGC

[BG76a] Alfred Brauer and Ivey C. Gentry. Bounds for the greatest characteristic root of an irreducible nonnegative matrix II. Linear Algebra and its Applications, 13(1–2):109–114, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379576900483.

Brualdi:1976:CPD

[BG76b] Richard A. Brualdi and Peter M. Gibson. Convex polyhedra of doubly stochastic matrices — IV. Linear Algebra and its Applications, 15(2):153–172, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379576900136.

Brawley:1978:IMF

[BG78] J. V. Brawley and R. O. Gamble. Involutory matrices over finite commutative rings. Linear Algebra and its Applications, 21(2):175– 188, August 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379578900411.

Brualdi:1979:UTN

[BH79] Richard A. Brualdi and Mark B. Hedrick. A unified treatment of nearly reducible and nearly decomposable matrices. Linear Algebra and its Applications, 24(??):51–73, April 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579901460.

Bellissard:1978:HFH

[BIL78] Jean Bellissard, Bruno Iochum, and Ricardo Lima. Homogeneous and facially homogeneous self-dual cones. Linear Algebra and its Applications, 19(1):1–16, ???? 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379578900022.

Barnett:1978:ISC

[BJ78] S. Barnett and E. I. Jury. Inners and Schur complement. Linear Algebra and its Applications, 22(??):57–63, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900575.

Butler:1973:CPT

[BK73] William Butler and Irving Kaplansky. A curious property of threeby-three matrices and an application to orders. Linear Algebra and its Applications, 6(??):269–272, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379573900281.

Berman:1978:LM

[BK78] Abraham Berman and Anton Kotzig. The length of a (0, 1) matrix. Linear Algebra and its Applications, 20(3):197–203, June 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900186.

Baksalary:1979:ME

[BK79a] J. K. Baksalary and R. Kala. The matrix equation AX − Y B = C. Linear Algebra and its Applications, 25(??):41–43, June 1979.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579900041.

Baksalary:1979:TRB

[BK79b] J. K. Baksalary and R. Kala. Two relations between oblique and Λorthogonal projectors. Linear Algebra and its Applications, 24(??):99– 103, April 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379579901502.

Barker:1975:SCM

[BL75] George Phillip Barker and Raphael Loewy. The structure of cones of matrices. Linear Algebra and its Applications, 12(1):87–94, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575901305.

Brualdi:1978:TAP

[BL78] Richard A. Brualdi and Geoffrey M. Lee. On the truncated assignment polytope. Linear Algebra and its Applications, 19(1):33–62, ???? 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900046.

Beard:1976:MFI

[BM76] Jacob T. B. Beard, Jr. and Robert M. McConnel. Matrix fields over the integers modulo m. Linear Algebra and its Applications, 14(2): 95–105, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379576900161.

Bhatia:1979:RCS

[BM79] Rajendra Bhatia and Kalyan K. Mukherjea. On the rate of change of spectra of operators. Linear Algebra and its Applications, 27(??):147– 157, October 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379579900375.

Bode:1979:SAR

[Bod79] Albert Bode. Symmetrien und antisymmetrien der Robert’schen Quasiordnung in matrizenräumen. (German) [Symmetries and antisymmetries in Robert’s quasiorder in matrix spaces]. Linear Algebra and its Applications, 23(??):87–99, February 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579900958.

Bose:1975:MFR

[Bos75] N. K. Bose. Matrix factorization in a real field. Linear Algebra and its Applications, 11(1):21–25, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379575901147.

Botta:1978:LMP

[Bot78] Peter Botta. Linear maps that preserve singular and nonsingular matrices. Linear Algebra and its Applications, 20(1):45–49, March 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900277.

Berman:1976:ETM

[BP76] A. Berman and R. J. Plemmons. Eight types of matrix monotonicity. Linear Algebra and its Applications, 13(1–2):115–123, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900495.

Brawley:1970:RBV

[Bra70] Joel Brawley, Jr. On the ranks of bases of vector spaces of matrices. Linear Algebra and its Applications, 3(1):51–55, January 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379570900285.

Brawley:1979:NPM

[Bra79] J. V. Brawley. A note on polynomial matrix functions over a finite field. Linear Algebra and its Applications, 28(??):35–38, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901150.

Brezinski:1974:SRT

[Bre74] Claude Brezinski. Some results in the theory of the vector ϵalgorithm. Linear Algebra and its Applications, 8(1):77–86, February 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957490010X.

Brezinski:1975:CEM

[Bre75] Claude Brezinski. Computation of the eigenelements of a matrix by the ϵ-algorithm. Linear Algebra and its Applications, 11(1):7–20, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575901135.

Brenner:1976:CAD

[Bre76] J. L. Brenner. Corrections to “Applications of the Dieudonné determinant”. Linear Algebra and its Applications, 13(3):289, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576901026.

Bridges:1978:ADT

[Bri78] W. G. Bridges. Algebraic duality theorems with combinatorial applications. Linear Algebra and its Applications, 22(??):157–162, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379578900666.

Barker:1975:APF

[BS75] G. P. Barker and Hans Schneider. Algebraic Perron–Frobenius theory. Linear Algebra and its Applications, 11(3):219–233, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900221.

Barker:1973:SOS

[BT73] G. P. Barker and R. E. L. Turner. Some observations on the spectra of cone preserving maps. Linear Algebra and its Applications, 6(??):149– 153, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/002437957390013X.

Butler:1972:NIM

[But72] Kim Ki-Hang Butler. The number of idempotents in (0,1)-matrix semigroups. Linear Algebra and its Applications, 5(3):233–246, July 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379572900055.

Berman:1978:PMN

[BVW78] A. Berman, R. S. Varga, and R. C. Ward. A L P S: Matrices with nonpositive off-diagonal entries. Linear Algebra and its Applications, 21(3):233–244, September 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/002437957890085X.

Botta:1970:USS

[BW70] Peter Botta and S. G. Williamson. Unitary similarity of symmetry operators. Linear Algebra and its Applications, 3(2):249–256, April 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379570900170.

Bruno:1971:PMM

[BW71] John Bruno and Louis Weinberg. The principal minors of a matroid. Linear Algebra and its Applications, 4(1):17–54, January 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379571900267.

Berman:1978:ACS

[BW78] Abraham Berman and Robert C. Ward. ALPS: Classes of stable and semipositive matrices. Linear Algebra and its Applications, 21(2):163– 174, August 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/002437957890040X.

Brenner:1979:PPD

[BW79] J. L. Brenner and Edward T. H. Wang. Permanental pairs of doubly stochastic matrices. II. Linear Algebra and its Applications, 28 (??):39–41, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379579901162.

Byers:1970:NAN

[Bye70] V. Byers. Non-Archimedian norms and bounds. Linear Algebra and its Applications, 3(1):57–77, January 1970. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379570900297.

Cain:1973:HCM

[Cai73] Bryan E. Cain. Hermitian congruence for matrices with semidefinite real part and for H-stable matrices. Linear Algebra and its Applications, 7(1):43–52, January 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379573900359.

Calderon:1973:NBF

[Cal73] A. P. Calderón. A note on biquadratic forms. Linear Algebra and its Applications, 7(2):175–177, April 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379573900529.

Campbell:1977:CMP

[Cam77] Stephen L. Campbell. On continuity of the Moore–Penrose and Drazin generalized inverses. Linear Algebra and its Applications, 18(1):53–57, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379577900799.

Campbell:1979:LBS

[Cam79] Stephen L. Campbell. Limit behavior of solutions of singular difference equations. Linear Algebra and its Applications, 23(??):167–178, February 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379579901009.

Carlson:1972:IDE

[Car72] David Carlson. Inequalities for the degrees of elementary divisors of modules. Linear Algebra and its Applications, 5(3):293–298, July 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379572900109.

Carlson:1976:GMM

[Car76] David Carlson. Generalizations of matrix monotonicity. Linear Algebra and its Applications, 13(1–2):125–131, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900501.

Cater:1978:PCC

[Cat78] F. S. Cater. Products of central collineations. Linear Algebra and its Applications, 19(3):251–274, ???? 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379578900149.

Cantoni:1976:EES

[CB76] A. Cantoni and P. Butler. Eigenvalues and eigenvectors of symmetric centrosymmetric matrices. Linear Algebra and its Applications, 13(3):275–288, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379576901014.

Carlson:1979:LME

[CD79] David H. Carlson and Biswa Nath Datta. The Lyapunov matrix equation SA + A^{∗}S = S^{∗}B^{∗}BS. Linear Algebra and its Applications, 28(??):43–52, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379579901174.

Cottle:1972:MTC

[CF72] Richard W. Cottle and Jacques A. Ferland. Matrix-theoretic criteria for the quasi-convexity and pseudo-convexity of quadratic functions. Linear Algebra and its Applications, 5(2):123–136, April 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379572900225.

Cline:1979:RDM

[CF79] R. E. Cline and R. E. Funderlic. The rank of a difference of matrices and associated generalized inverses. Linear Algebra and its Applications, 24(??):185–215, April 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579901587.

Charnes:1977:PAI

[CGG77] A. Charnes, Daniel Granot, and Frieda Granot. A primal algorithm for interval linear-programming problems. Linear Algebra and its Applications, 17(1):65–78, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379577900428.

Cheung:1975:ISP

[CH75] Shiu Ming Cheung and Harry Hochstadt. An inverse spectral problem. Linear Algebra and its Applications, 12(3):215–222, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900439.

Carlson:1976:GCI

[CH76] David Carlson and Richard Hill. Generalized controllability and inertia theory. Linear Algebra and its Applications, 15(2):177–187, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957690015X.

Carlitz:1977:EMG

[CH77] L. Carlitz and John H. Hodges. Enumeration of matrices of given rank with submatrices of given rank. Linear Algebra and its Applications, 16(3):285–291, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379577900106.

Chao:1970:RCM

[Cha70] Chong-Yun Chao. A remark on cyclic matrices. Linear Algebra and its Applications, 3(2):165–172, April 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379570900121.

Chao:1973:TC

[Cha73] Chong-Yun Chao. On a type of circulants. Linear Algebra and its Applications, 6(??):241–248, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379573900244.

Chakravarti:1975:CIN

[Cha75] I. M. Chakravarti. On a characterization of irreducibility of a nonnegative matrix. Linear Algebra and its Applications, 10(2):103– 109, April 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379575900026.

Chan:1979:CTS

[Cha79] Gin-Hor Hor Chan. (k)-characters and the triviality of symmetry classes. Linear Algebra and its Applications, 25(??):139–149, June 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579900132.

Ching:1977:LEC

[Chi77] Wai-Sin Ching. Linear equations over commutative rings. Linear Algebra and its Applications, 18(3):257–266, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900556.

Cottle:1970:CCM

[CHL70] R. W. Cottle, G. J. Habetler, and C. E. Lemke. On classes of copositive matrices. Linear Algebra and its Applications, 3(3):295–310, July 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379570900029.

Choi:1975:CPL

[Cho75a] Man-Duen Choi. Completely positive linear maps on complex matrices. Linear Algebra and its Applications, 10(3):285–290, June 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900750.

Choi:1975:PSB

[Cho75b] Man-Duen Choi. Positive semidefinite biquadratic forms. Linear Algebra and its Applications, 12(2):95–100, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900580.

Christian:1979:MID

[Chr79] Floyd L. Christian, Jr. A k-measure of irreducibility and doubly stochastic matrices. Linear Algebra and its Applications, 24(??):225– 237, April 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379579901605.

Ciampi:1970:CCM

[Cia70] Rita Ciampi. Characterization of a class of matrices as sums of squares. Linear Algebra and its Applications, 3(1):45–50, January 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379570900273.

Carlson:1974:RLT

[CL74] David Carlson and Raphael Loewy. On ranges of Lyapunov transformations. Linear Algebra and its Applications, 8(3):237–248, June 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957490069X.

Claasen:1978:MCL

[Cla78a] H. L. Claasen. The multiplications in GF(q)[x]/(a(x)) considered as linear transformations. Linear Algebra and its Applications, 22 (??):105–123, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379578900629.

Claessens:1978:LZG

[Cla78b] Guido Claessens. On the location of the zeros of a generalized polynomial. Linear Algebra and its Applications, 22(??):79–88, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900599.

Campbell:1975:CPD

[CM75] Stephen L. Campbell and Carl D. Meyer, Jr. Continuity properties of the Drazin pseudoinverse. Linear Algebra and its Applications, 10(1):77–83, February 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/002437957590097X.

Campbell:1978:WDI

[CM78] Stephen L. Campbell and Carl D. Meyer, Jr. Weak Drazin inverses. Linear Algebra and its Applications, 20(2):167–178, ???? 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900484.

Cohen:1974:NPS

[Coh74] A. M. Cohen. A note on pivot size in Gaussian elimination. Linear Algebra and its Applications, 8(4):361–368, August 1974.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379574900640.

Cohn:1978:TFI

[Coh78] Harvey Cohn. Ternary forms as invariants of Markoff forms and other SL _{2} (Z)-bundles. Linear Algebra and its Applications, 21(1): 3–12, July 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379587901960.

Cook:1975:EIM

[Coo75] P. A. Cook. Estimates for the inverse of a matrix. Linear Algebra and its Applications, 10(1):41–53, February 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379575900932.

Cordes:1975:NPP

[Cor75] Craig M. Cordes. A note on Pall partitions over finite fields. Linear Algebra and its Applications, 12(1):81–85, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575901299.

Cottle:1974:MSC

[Cot74] Richard W. Cottle. Manifestations of the Schur complement. Linear Algebra and its Applications, 8(3):189–211, June 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379574900664.

Courter:1973:MCS

[Cou73] R. C. Courter. Maximal commutative subalgebras of K _{n} at exponent three. Linear Algebra and its Applications, 6(??):1–11, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379573900025.

Calogero:1979:SDR

[CP79] F. Calogero and A. M. Perelomov. Some Diophantine relations involving circular functions of rational angles. Linear Algebra and its Applications, 25(??):91–94, June 1979. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579900089.

Cline:1974:GIC

[CPW74] R. E. Cline, R. J. Plemmons, and G. Worm. Generalized inverses of certain Toeplitz matrices. Linear Algebra and its Applications, 8(1):25–33, February 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379574900044.

Cross:1978:TTM

[Cro78] G. W. Cross. Three types of matrix stability. Linear Algebra and its Applications, 20(3):253–263, June 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379578900216.

Cruse:1975:PFC

[Cru75] Allan B. Cruse. A proof of Fulkerson’s characterization of permutation matrices. Linear Algebra and its Applications, 12(1):21–28, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957590124X.

Cruse:1977:SCP

[Cru77] Allan B. Cruse. Some combinatorial properties of centrosymmetric matrices. Linear Algebra and its Applications, 16(1):65–77, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900209.

Cruse:1979:RVA

[Cru79] Allan B. Cruse. On removing a vertex from the assignment polytope. Linear Algebra and its Applications, 26(??):45–57, August 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957990171X.

Cryer:1973:FTP

[Cry73] Colin W. Cryer. The LU-factorization of totally positive matrices. Linear Algebra and its Applications, 7(1):83–92, January 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379573900396.

Cryer:1976:SPT

[Cry76] Colin W. Cryer. Some properties of totally positive matrices. Linear Algebra and its Applications, 15(1):1–25, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900768.

Cunningham:1977:UMI

[Cun77] W. H. Cunningham. An unbounded matroid intersection polyhedron. Linear Algebra and its Applications, 16(3):209–215, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900039.

Carlson:1973:MF

[CV73a] David H. Carlson and Richard S. Varga. Minimal G-functions. Linear Algebra and its Applications, 6(??):97–117, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379573900098.

Carlson:1973:MFI

[CV73b] David H. Carlson and Richard S. Varga. Minimal G-functions. II. Linear Algebra and its Applications, 7(3):233–242, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379573900426.

Carlson:1974:CF

[CV74] David H. Carlson and Richard S. Varga. On collections of Gfunctions. Linear Algebra and its Applications, 8(1):65–76, February 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379574900093.

Datta:1978:EPS

[Dat78a] B. N. Datta. An elementary proof of the stability criterion of Liénard and Chipart. Linear Algebra and its Applications, 22(??):89–96, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379578900605.

Datta:1978:RHF

[Dat78b] B. N. Datta. On the Routh–Hurwitz–Fujiwara and the Schur– Cohn–Fujiwara theorems for the root-separation problem. Linear Algebra and its Applications, 22(??):235–245, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900745.

Datta:1978:SS

[Dat78c] Biswa Nath Datta. Stability and D-stability. Linear Algebra and its Applications, 21(2):135–141, August 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437957890037X.

Daughtry:1978:ISQ

[Dau78] John Daughtry. Isolated solutions of quadratic matrix equations. Linear Algebra and its Applications, 21(1):89–94, July 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379587902023.

Davis:1970:RCP

[Dav70] Chandler Davis. Representing a commuting pair by tensor products. Linear Algebra and its Applications, 3(3):355–357, July 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379570900054.

Davis:1973:EFC

[Dav73] Chandler Davis. Explicit functional calculus. Linear Algebra and its Applications, 6(??):193–199, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379573900190.

Davis:1976:EPE

[Dav76] Chandler Davis. An extremal problem for extensions of a sesquillinear form. Linear Algebra and its Applications, 13(1–2):91–102, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957690046X.

Davis:1977:GAD

[Dav77] Chandler Davis. Geometric approach to a dilation theorem. Linear Algebra and its Applications, 18(1):33–43, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900775.

Davis:1979:CTG

[Dav79] Philip J. Davis. Cyclic transformations of n-gons and related quadratic forms. Linear Algebra and its Applications, 25(??):57– 75, June 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379579900065.

DePillis:1971:GES

[DB71] John De Pillis and Joel Brenner. Generalized elementary symmetric functions and quaternion matrices. Linear Algebra and its Applications, 4(1):55–69, January 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379571900279.

deBoor:1978:NSR

[dBG78] C. de Boor and G. H. Golub. The numerically stable reconstruction of a Jacobi matrix from spectral data. Linear Algebra and its Applications, 21(3):245–260, September 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379578900861.

Doob:1979:SCE

[DC79] Michael Doob and Dragoš Cvetković. On spectral characterizations and embeddings of graphs. Linear Algebra and its Applications, 27(??):17–26, October 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379579900284.

Della-Dora:1975:NLA

[DD75] J. Della-Dora. Numerical linear algorithms and group theory. Linear Algebra and its Applications, 10(3):267–283, June 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900749.

Datta:1976:ACP

[DD76] B. N. Datta and Karabi Datta. An algorithm for computing powers of a Hessenberg matrix and its applications. Linear Algebra and its Applications, 14(3):273–284, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379576900720.

DePillis:1971:IPP

[De 71] John De Pillis. Inequalities for partitioned positive semidefinite matrices. Linear Algebra and its Applications, 4(1):79–94, January 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379571900309.

Deutsch:1970:MNZ

[Deu70] Emeric Deutsch. Matricial norms and the zeros of polynomials. Linear Algebra and its Applications, 3(4):483–489, October 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379570900388.

Deutsch:1971:SIR

[Deu71] Emeric Deutsch. Semi-inverses, reflexive semi-inverses, and pseudoinverses of an arbitrary linear transformation. Linear Algebra and its Applications, 4(4):313–322, October 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379571900024.

Deutsch:1973:MNZ

[Deu73] Emeric Deutsch. Matricial norms and the zeros of lacunary polynomials. Linear Algebra and its Applications, 6(??):143–148, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379573900128.

Deutsch:1978:SMP

[Deu78] Emeric Deutsch. On the spread of matrices and polynomials. Linear Algebra and its Applications, 22(??):49–55, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900563.

Deddens:1975:RLT

[DF75] J. A. Deddens and P. A. Fillmore. Reflexive linear transformations. Linear Algebra and its Applications, 10(1):89–93, February 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900993.

Deutsch:1976:FPT

[DGS76] Emeric Deutsch, P. M. Gibson, and Hans Schneider. The Fuglede– Putnam theorem and normal products of matrices. Linear Algebra and its Applications, 13(1–2):53–58, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379576900422.

DiasdaSilva:1979:CED

[Dia79] J. A. Dias da Silva. Conditions for equality of decomposable symmetric tensors. Linear Algebra and its Applications, 24(??):85–92, April 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901484.

Dixon:1976:NLN

[Dix76] John D. Dixon. Nilpotence and local nilpotence of linear groups. Linear Algebra and its Applications, 13(1–2):59–67, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900434.

Djokovic:1973:SBL

[Djo73] Dragomir Ž. Djoković. On the similarity between a linear transformation and its adjoint. Linear Algebra and its Applications, 6(??):155– 158, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379573900141.

Djokovic:1975:CHS

[Djo75] Dragomir Ž. Djoković. Characterization of Hermitian and skew-Hermitian maps between matrix algebras. Linear Algebra and its Applications, 12(2):165–170, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379575900646.

Djokovic:1976:PSD

[Djo76] Dragomir Ž. Djoković. Positive semi-definite matrices as sums of squares. Linear Algebra and its Applications, 14(1):37–40, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900616.

Djokovic:1978:CPC

[Djo78] Dragomir Ž. Djoković. Classification of pairs consisting of a linear and a semilinear map. Linear Algebra and its Applications, 20(2):147– 165, ???? 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379578900472.

Deutsch:1978:MLD

[DM78] Emeric Deutsch and Max Mlynarski. Matricial logarithmic derivatives. Linear Algebra and its Applications, 19(1):17–31, ???? 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900034.

Djokovic:1979:PRG

[DM79] Dragomir Ž. Djoković and Jerry Malzan. Products of reflections in the general linear group over a division ring. Linear Algebra and its Applications, 28(??):53–62, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579901186.

deOliveira:1972:CVM

[dO72] G. N. de Oliveira. On the characteristic vectors of a matrix. Linear Algebra and its Applications, 5(3):189–196, July 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379572900018.

deOliveira:1974:NCR

[dO74] G. N. de Oliveira. Note on the characteristic roots of tournament matrices. Linear Algebra and its Applications, 8(3):271–272, June 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957490072X.

deOliveira:1979:CED

[dOdS79] G. N. de Oliveira and J. A. Dias da Silva. Conditions for equality of decomposable symmetric tensors. II. Linear Algebra and its Applications, 28(??):161–176, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579901307.

Donnelly:1971:PCR

[Don71] J. D. P. Donnelly. Periodic chaotic relaxation. Linear Algebra and its Applications, 4(2):117–128, April 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379571900334.

Doob:1970:CCG

[Doo70] Michael Doob. On characterizing certain graphs with four eigenvalues by their spectra. Linear Algebra and its Applications, 3(4):461–482, October 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379570900376.

Doob:1975:SCL

[Doo75] Michael Doob. A spectral characterization of the line graph of a BIBD with λ = 1. Linear Algebra and its Applications, 12(1):11– 20, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379575901238.

Downs:1971:IMR

[Dow71] T. Downs. On the inversion of a matrix of rational functions. Linear Algebra and its Applications, 4(1):1–10, January 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379571900243.

Dunkl:1976:LPP

[DR76] Charles F. Dunkl and Donald E. Ramirez. A linear programming problem in harmonic analysis. Linear Algebra and its Applications, 14(2):107–116, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379576900173.

Dlab:1977:NFR

[DR77] Vlastimil Dlab and Claus Michael Ringel. Normal forms of real matrices with respect to complex similarity. Linear Algebra and its Applications, 17(2):107–124, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379577900313.

Dekker:1971:SAH

[DT71] T. J. Dekker and J. F. Traub. The shifted QR algorithm for Hermitian matrices. Linear Algebra and its Applications, 4(2):137–154, April 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379571900358.

Duke:1970:NEL

[Duk70] John W. Duke. A note on EP linear transformations. Linear Algebra and its Applications, 3(3):379–382, July 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379570900078.

Eaves:1973:PLR

[Eav73] B. Curtis Eaves. Piecewise linear retractions by reflexion. Linear Algebra and its Applications, 7(2):93–98, April 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379573900463.

Eckhardt:1975:TDC

[Eck75] Ulrich Eckhardt. Theorems on the dimension of convex sets. Linear Algebra and its Applications, 12(1):63–76, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575901275.

Eisenfeld:1973:BDP

[Eis73] J. Eisenfeld. Block diagonalization of partitioned matrix operators. Linear Algebra and its Applications, 6(??):183–191, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379573900189.

Eisenfeld:1976:BDE

[Eis76] Jerome Eisenfeld. Block diagonalization and eigenvalues. Linear Algebra and its Applications, 15(3):205–215, ???? 1976.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900264.

Kuo:1973:SPT

[eK73] Yueh er Kuo. Some properties of tensor products of matrices. Linear Algebra and its Applications, 7(1):63–70, January 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379573900372.

Ellers:1979:PTI

[Ell79] Erich W. Ellers. Products of two involutory matrices over skewfields. Linear Algebra and its Applications, 26(??):59–63, August 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901721.

Elsner:1974:QUH

[Els74] Ludwig Elsner. Quasimonotonie und Ungleichungen in halbgeordneten Räumen. (German) [Quasimonotonicity and inequalities in partially-ordered spaces]. Linear Algebra and its Applications, 8 (3):249–261, June 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379574900706.

Elsner:1976:NCI

[Els76a] L. Elsner. A note on characterizations of irreducibility of nonnegative matrices. Linear Algebra and its Applications, 14(2):187–188, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900239.

Elsner:1976:IIC

[Els76b] Ludwig Elsner. Inverse iteration for calculating the spectral radius of a non-negative irreducible matrix. Linear Algebra and its Applications, 15(3):235–242, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/002437957690029X.

Elsner:1979:SAP

[Els79] L. Elsner. On some algebraic problems in connection with general eigenvalue algorithms. Linear Algebra and its Applications, 26 (??):123–138, August 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379579901757.

Engel:1973:RES

[Eng73] Gernot M. Engel. Regular equimodular sets of matrices for generalized matrix functions. Linear Algebra and its Applications, 7(3):243– 274, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379573900438.

Engel:1973:CDP

[ES73] Gernot M. Engel and Hans Schneider. Cyclic and diagonal products on a matrix. Linear Algebra and its Applications, 7(4):301–335, October 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0024379573800035.

Estes:1979:DGA

[Est79] Dennis R. Estes. Determinants of Galois automorphisms of maximal commutative rings of 2 × 2 matrices. Linear Algebra and its Applications, 27(??):225–243, October 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579900454.

Fan:1972:GCT

[Fan72] Ky Fan. Generalized Cayley transforms and strictly dissipative matrices. Linear Algebra and its Applications, 5(2):155–172, April 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379572900250.

Fan:1975:TAC

[Fan75] Ky Fan. Two applications of a consistency theorem for systems of linear inequalities. Linear Algebra and its Applications, 11(2):171– 180, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379575900567.

Farjot:1976:RMH

[Far76] Paulette Farjot. Rang des matrices Hermitiennes semi-définies positives indéfiniment divisibles. (French) [Rank of semidefinite positive infinitely-divisible Hermitian matrices]. Linear Algebra and its Applications, 15(3):189–196, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379576900240.

Farebrother:1977:NSC

[Far77] R. W. Farebrother. Necessary and sufficient conditions for a quadratic form to be positive whenever a set of homogeneous linear constraints is satisfied. Linear Algebra and its Applications, 16(1):39–42, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900179.

Feintuch:1977:ODP

[Fei77] Avraham Feintuch. One-dimensional perturbations of cyclic operators. Linear Algebra and its Applications, 18(2):171–174, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900726.

Ferland:1978:MCP

[Fer78a] Jacques A. Ferland. Matrix criteria for pseudoconvex functions in the class C^{2}. Linear Algebra and its Applications, 21(1):47–57, July 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379587901996.

Feroe:1978:SRT

[Fer78b] John A. Feroe. A spectral radius theorem for matrix seminorms. Linear Algebra and its Applications, 22(??):97–103, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900617.

Fitz-Gerald:1977:CMG

[FG77] D. G. Fitz-Gerald. Computing the maximum generalized inverse of a Boolean matrix. Linear Algebra and its Applications, 16(3):203– 207, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379577900027.

Fillmore:1977:HSL

[FHL77] P. A. Fillmore, Domingo A. Herrero, and W. E. Longstaff. The hyperinvariant subspace lattice of a linear transformation. Linear Algebra and its Applications, 17(2):125–132, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379577900325.

Fiedler:1972:BED

[Fie72] Miroslav Fiedler. Bounds for eigenvalues of doubly stochastic matrices. Linear Algebra and its Applications, 5(3):299–310, July 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379572900110.

Fenner:1974:MBS

[FL74a] T. I. Fenner and G. Loizou. Matrix Bounds on the Spectral Condition Number. Linear Algebra and its Applications, 8(2):157–178, April 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379574900536.

Freedman:1974:SCE

[FL74b] H. I. Freedman and J. D. Lawson. Systems with constant eigenvectors with applications to exact and numerical solutions of ordinary differential equations. Linear Algebra and its Applications, 8(4):369–374, August 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379574900652.

Flanders:1974:NPR

[Fla74] Harley Flanders. A new proof of R. Foster’s averaging formula in networks. Linear Algebra and its Applications, 8(1):35–37, February 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379574900056.

Fletcher:1976:FSI

[Fle76] R. Fletcher. Factorizing symmetric indefinite matrices. Linear Algebra and its Applications, 14(3):257–272, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379576900719.

FloresDeChela:1979:GIN

[Flo79] Delia Flores De Chela. Generalized inverses on normed vector spaces. Linear Algebra and its Applications, 26(??):243–263, August 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901800.

Friedland:1978:BSD

[FM78] Shmuel Friedland and Charles A. Micchelli. Bounds on the solutions of difference equations and spline interpolation at knots. Linear Algebra and its Applications, 20(3):219–251, June 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900204.

Friedland:1979:ENJ

[FM79] S. Friedland and A. A. Melkman. On the eigenvalues of nonnegative Jacobi matrices. Linear Algebra and its Applications, 25(??):239– 253, June 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379579900211.

Friedlander:1979:NIF

[FMKL79] B. Friedlander, M. Morf, T. Kailath, and L. Ljung. New inversion formulas for matrices classified in terms of their distance from Toeplitz matrices. Linear Algebra and its Applications, 27(??):31–60, October 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579900302.

Fixman:1978:EPL

[FO78] Uri Fixman and Frank Okoh. Extensions of pairs of linear transformations between infinite-dimensional vector spaces. Linear Algebra and its Applications, 19(3):275–291, ???? 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379578900150.

Fox:1978:CRT

[Fox78] David W. Fox. The convexity of the range of three Hermitian forms and of the numerical range of sesquilinear forms. Linear Algebra and its Applications, 22(??):191–194, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379578900708.

Freese:1973:IGM

[Fre73] Ralph Freese. Inequalities for generalized matrix functions based on arbitrary characters. Linear Algebra and its Applications, 7(4):337– 345, October 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0024379573800047.

Friedland:1975:IME

[Fri75] Shmuel Friedland. On inverse multiplicative eigenvalue problems for matrices. Linear Algebra and its Applications, 12(2):127–137, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900610.

Friedland:1977:IEP

[Fri77] Shmuel Friedland. Inverse eigenvalue problems. Linear Algebra and its Applications, 17(1):15–51, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379577900398.

Friedland:1979:CTA

[Fri79] Shmuel Friedland. A characterization of transform absolute norms. Linear Algebra and its Applications, 28(??):63–68, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901198.

Felzenbaum:1979:CPN

[FT79] A. Felzenbaum and A. Tamir. Cyclic permutations and nearly symmetric integer vectors. Linear Algebra and its Applications, 27 (??):159–166, October 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379579900387.

Fu:1975:ITJ

[Fu75] Lorraine Fu. An inverse theorem for Jacobi matrices. Linear Algebra and its Applications, 12(1):55–61, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379575901263.

Fulton:1973:CEL

[Ful73a] John D. Fulton. Characterization and enumeration of linear classes of involutions over a finite field. Linear Algebra and its Applications, 6(??):119–127, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379573900104.

Fulton:1973:LCI

[Ful73b] John D. Fulton. Linear classes of involutions over fields of characteristic two. Linear Algebra and its Applications, 6(??):129–142, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379573900116.

Fulton:1979:GIM

[Ful79] John D. Fulton. Generalized inverses of matrices over fields of characteristic two. Linear Algebra and its Applications, 28(??):69–76, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379579901204.

Funderlic:1979:SCO

[Fun79] R. E. Funderlic. Some characterizations of orthant monotonic norms. Linear Algebra and its Applications, 28(??):77–83, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901216.

Flanders:1974:POP

[FW74] Harley Flanders and Harald K. Wimmer. Positive operators and a problem in control theory. Linear Algebra and its Applications, 8(1):39–42, February 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379574900068.

Gallai:1978:LSW

[Gal78] Tibor Gallai. The life and scientific work of Dénes König (1884– 1944). Linear Algebra and its Applications, 21(3):189–205, September 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900824.

Gentleman:1975:EAD

[Gen75] W. M. Gentleman. Error analysis of QR decompositions by Givens transformations. Linear Algebra and its Applications, 10(3):189– 197, June 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379575900683.

Gerstein:1977:LAM

[Ger77] Larry J. Gerstein. A local approach to matrix equivalence. Linear Algebra and its Applications, 16(3):221–232, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900052.

Gottlieb:1977:MEA

[GG77] David Gottlieb and Max D. Gunzburger. On the matrix equations AH + HA^{∗} = A^{∗}H + HA = I. Linear Algebra and its Applications, 17(3):277–282, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379577900659.

Gustafson:1976:PI

[GHR76] W. H. Gustafson, P. R. Halmos, and H. Radjavi. Products of involutions. Linear Algebra and its Applications, 13(1–2):157–162, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900549.

Gibson:1976:PBD

[Gib76] Peter M. Gibson. Products of basic doubly stochastic matrices over a field. Linear Algebra and its Applications, 15(2):99–118, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900112.

Gibson:1978:RPR

[Gib78] Peter M. Gibson. Real permanental roots of doubly stochastic matrices. Linear Algebra and its Applications, 21(3):289–291, September 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900897.

Gilman:1977:MRA

[Gil77] Jane Gilman. A matrix representation for automorphisms of compact Riemann surfaces. Linear Algebra and its Applications, 17(2):139– 147, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379577900349.

Gastinel:1970:CNG

[GJ70] N. Gastinel and J. L. Joly. Condition numbers and general projection method. Linear Algebra and its Applications, 3(2):185–224, April 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379570900157.

Gorelick:1976:EKS

[GK76] Michael R. Gorelick and Herbert C. Kranzer. An extension of the Kreiss stability theorem to families of matrices of unbounded order. Linear Algebra and its Applications, 14(3):237–255, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900707.

Gohberg:1978:SAMa

[GLR78a] I. Gohberg, P. Lancaster, and L. Rodman. Spectral analysis of matrix polynomials — I. Canonical forms and divisors. Linear Algebra and its Applications, 20(1):1–44, March 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379578900265.

Gohberg:1978:SAMb

[GLR78b] I. Gohberg, P. Lancaster, and L. Rodman. Spectral analysis of matrix polynomials — II. The resolvent form and spectral divisors. Linear Algebra and its Applications, 21(1):65–88, July 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379587902011.

Gill:1973:NSF

[GM73] Philip E. Gill and Walter Murray. A numerically stable form of the simplex algorithm. Linear Algebra and its Applications, 7(2):99– 138, April 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379573900475.

Giles:1979:TDI

[GP79] F. R. Giles and W. R. Pulleyblank. Total dual integrality and integer polyhedra. Linear Algebra and its Applications, 25(??):191–196, June 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579900181.

Gohberg:1979:SSM

[GR79] I. Gohberg and L. Rodman. On the spectral structure of monic matrix polynomials and the extension problem. Linear Algebra and its Applications, 24(??):157–172, April 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579901563.

Graver:1975:BDS

[Gra75] Jack E. Graver. Boolean designs and self-dual matroids. Linear Algebra and its Applications, 10(2):111–128, April 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900038.

Greville:1973:SFA

[Gre73] T. N. E. Greville. The Souriau–Frame algorithm and the Drazin pseudoinverse. Linear Algebra and its Applications, 6(??):205–208, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379573900219.

Green:1974:WSS

[Gre74] B. M. Green. On weakly sign-symmetric matrices. Linear Algebra and its Applications, 8(4):355–360, August 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379574900639.

Grone:1977:NDO

[Gro77] Robert Grone. A note on the dimension of an orbital subspace. Linear Algebra and its Applications, 17(3):283–286, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900660.

Goldberg:1976:IRI

[GS76] Moshe Goldberg and Ernst G. Straus. Inclusion relations involving k-numerical ranges. Linear Algebra and its Applications, 15(3):261– 270, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/002437957690032X.

Goldberg:1977:EIR

[GS77] Moshe Goldberg and E. G. Straus. Elementary inclusion relations for generalized numerical ranges. Linear Algebra and its Applications, 18(1):1–24, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379577900751.

Goldberg:1979:NPN

[GS79] Moshe Goldberg and E. G. Straus. Norm properties of C-numerical radii. Linear Algebra and its Applications, 24(??):113–131, April 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901526.

Greville:1979:BMT

[GT79] T. N. E. Greville and W. F. Trench. Band matrices with Toeplitz inverses. Linear Algebra and its Applications, 27(??):199–209, October 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579900429.

Goldberg:1975:NRP

[GTZ75] Moshe Goldberg, Eitan Tadmor, and Gideon Zwas. Numerical radius of positive matrices. Linear Algebra and its Applications, 12(3):209– 214, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379575900427.

Gustafson:1979:RTC

[Gus79] William H. Gustafson. Roth’s theorems over commutative rings. Linear Algebra and its Applications, 23(??):245–251, February 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957990106X.

Golub:1979:UPD

[GV79] Gene H. Golub and Charles F. Van Loan. Unsymmetric positive definite linear systems. Linear Algebra and its Applications, 28(??):85–97, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379579901228.

Gray:1976:CJM

[GW76] L. J. Gray and D. G. Wilson. Construction of a Jacobi matrix from spectral data. Linear Algebra and its Applications, 14(2):131–134, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379576900203.

Garcia:1979:UM

[GZ79a] C. B. García and W. I. Zangwill. On univalence and P -matrices. Linear Algebra and its Applications, 24(??):239–250, April 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901617.

Gustafson:1979:MEM

[GZ79b] William H. Gustafson and Julius M. Zelmanowitz. On matrix equivalence and matrix equations. Linear Algebra and its Applications, 27(??):219–224, October 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379579900442.

Halmos:1971:EA

[Hal71] P. R. Halmos. Eigenvectors and adjoints. Linear Algebra and its Applications, 4(1):11–15, January 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379571900255.

Hald:1976:IEP

[Hal76a] Ole H. Hald. Inverse eigenvalue problems for Jacobi matrices. Linear Algebra and its Applications, 14(1):63–85, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900641.

Hall:1976:IBG

[Hal76b] Frank J. Hall. On the independence of blocks of generalized inverses of bordered matrices. Linear Algebra and its Applications, 14(1):

53–61, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/002437957690063X.

Hald:1979:NSG

[Hal79] Ole H. Hald. Numerical solution of the Gel’fand–Levitan equation. Linear Algebra and its Applications, 28(??):99–111, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957990123X.

Hartfiel:1975:TTG

[Har75a] D. J. Hartfiel. Two theorems generalizing the mean transition probability results in the theory of Markov chains. Linear Algebra and its Applications, 11(2):181–187, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379575900579.

Hartwig:1975:IIB

[Har75b] Robert E. Hartwig. 1–2 Inverses and the invariance of BA^{+}C. Linear Algebra and its Applications, 11(3):271–275, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900269.

Hartfiel:1977:MCN

[Har77] D. J. Hartfiel. Mathematical considerations of nonhomogeneous production processes. Linear Algebra and its Applications, 16(2):95– 107, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379577900246.

Hartwig:1978:SIR

[Har78] Robert E. Hartwig. Spectral inverses and the row-space equations. Linear Algebra and its Applications, 20(1):57–68, March 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900290.

Haynsworth:1970:RMU

[Hay70] Emilie V. Haynsworth. Reduction of a matrix using properties of the Schur complement. Linear Algebra and its Applications, 3(1):23–29, January 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/002437957090025X.

Hearon:1970:RIC

[Hea70] John Z. Hearon. Roots of an irreducible continuant. Linear Algebra and its Applications, 3(1):125–127, January 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379570900327.

Heading:1973:GZJ

[Hea73] J. Heading. Generalized Zilch and the Jordan canonical form. Linear Algebra and its Applications, 6(??):19–36, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379573900049.

Heading:1975:CGA

[Hea75] John Heading. Construction of generalized anisotropic zilch. Linear Algebra and its Applications, 11(3):277–289, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900270.

Hearon:1977:NS

[Hea77] John Z. Hearon. Nonsingular solutions of T A−BT = C. Linear Algebra and its Applications, 16(1):57–63, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379577900192.

Hedrick:1975:MDS

[Hed75] Mark Blondeau Hedrick. p-Minors of a doubly stochastic matrix at which the permanent achieves a minimum. Linear Algebra and its Applications, 10(2):177–179, April 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379575900099.

Heinz:1978:TPO

[Hei78] Tony F. Heinz. Topological properties of orthostochastic matrices. Linear Algebra and its Applications, 20(3):265–269, June 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900228.

Heuvers:1973:TFW

[Heu73] Konrad John Heuvers. On the types of functions which can serve as scalar products in a complex linear space. Linear Algebra and its Applications, 6(??):83–96, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379573900086.

Haynsworth:1976:EOP

[HFP76] Emilie Haynsworth, Miroslav Fiedler, and Vlastimil Pták. Extreme operators on polyhedral cones. Linear Algebra and its Applications, 13(1–2):163–172, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379576900550.

Hill:1973:LTW

[Hil73] Richard D. Hill. Linear transformations which preserve Hermitian matrices. Linear Algebra and its Applications, 6(??):257–262, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379573900268.

Hill:1977:ELU

[Hil77] Richard D. Hill. Eigenvalue location using certain matrix functions and geometric curves. Linear Algebra and its Applications, 16(1): 83–91, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379577900222.

Hills:1978:TMS

[Hil78] Norman L. Hills. Transposing a matrix by similarity operations. Linear Algebra and its Applications, 22(??):163–166, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900678.

Hirzebruch:1974:PPE

[Hir74] U. Hirzebruch. Produkte positiver elemente in formal-reellen Jordan-Algebren. (German) [Products of positive elements in formal-real Jordan algebras]. Linear Algebra and its Applications, 8(3):263–269, June 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379574900718.

Hartwig:1976:PEE

[HK76] Robert E. Hartwig and Irving J. Katz. Products of EP elements in reflexive semigroups. Linear Algebra and its Applications, 14(1): 11–19, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379576900586.

Horne:1970:MEB

[HM70] G. J. Horne and L. Magagna. Monotonicity of the eigenvalues of birth and death matrices. Linear Algebra and its Applications, 3(1):31–43, January 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379570900261.

Hung:1975:MPI

[HM75] Ching-hsiang Hung and T. L. Markham. The Moore–Penrose inverse of a partitioned matrix M = (AD; BC). Linear Algebra and its Applications, 11(1):73–86, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379575901184.

Hoskins:1978:OIP

[HM78] W. D. Hoskins and D. S. Meek. Order invariant properties of some matrices. Linear Algebra and its Applications, 21(3):261–269, September 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900873.

Hoskins:1979:HOI

[HMW79] W. D. Hoskins, D. S. Meek, and D. J. Walton. High-order iterative methods for the solution of the matrix equation XA + AY = F. Linear Algebra and its Applications, 23(??):121–139, February 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579900971.

Hochstadt:1970:CAC

[Hoc70] Harry Hochstadt. A constructive approach to the classical canonical form of linear operators. Linear Algebra and its Applications, 3(3):347–353, July 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379570900042.

Hochstadt:1975:THM

[Hoc75] Harry Hochstadt. On the theory of Hill’s matrices and related inverse spectral problems. Linear Algebra and its Applications, 11(1):41– 52, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379575901160.

Hochstadt:1979:CJM

[Hoc79] Harry Hochstadt. On the construction of a Jacobi matrix from mixed given data. Linear Algebra and its Applications, 28(??):113–115, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379579901241.

Hoffman:1972:EPE

[Hof72] A. J. Hoffman. Eigenvalues and partitionings of the edges of a graph. Linear Algebra and its Applications, 5(2):137–146, April 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379572900237.

Hoffman:1976:TUC

[Hof76a] Alan J. Hoffman. Total unimodularity and combinatorial theorems. Linear Algebra and its Applications, 13(1–2):103–108, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900471.

Hoffman:1976:OUA

[Hof76b] Frederick Hoffman. Orbits under actions of affine groups over GF (2). Linear Algebra and its Applications, 13(3):173–176, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900938.

Hoffman:1977:GWL

[Hof77] A. J. Hoffman. On graphs whose least eigenvalue exceeds −1 − \sqrt{2}. Linear Algebra and its Applications, 16(2):153–165, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900271.

Holbrook:1977:DCC

[Hol77] John A. R. Holbrook. Distortion coefficients for cryptocontractions. Linear Algebra and its Applications, 18(3):229–256, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900544.

Holbrook:1978:DCC

[Hol78] John A. R. Holbrook. Distortion coefficients for crypto-unitary operators. Linear Algebra and its Applications, 19(3):189–205, ???? 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900113.

Holgate:1979:CMG

[Hol79] P. Holgate. Canonical multiplication in the genetic algebra for linked loci. Linear Algebra and its Applications, 26(??):281–287, August 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901824.

Horn:1975:IBH

[Hor75] Roger A. Horn. On inequalities between Hermitian and symmetric forms. Linear Algebra and its Applications, 11(3):189–218, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957590021X.

Horne:1978:ICP

[Hor78a] J. G. Horne. On ideals of cone preserving maps. Linear Algebra and its Applications, 21(2):95–109, August 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379578900344.

Horne:1978:AGC

[Hor78b] J. G. Horne. On the automorphism group of a cone. Linear Algebra and its Applications, 21(2):111–121, August 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900356.

Householder:1971:MZT

[Hou71a] A. S. Householder. Multigradients and the zeros of transcendental functions. Linear Algebra and its Applications, 4(2):175–182, April 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379571900383.

Householder:1971:PTF

[Hou71b] A. S. Householder. The Padé table, the Frobenius identities, and the qd algorithm. Linear Algebra and its Applications, 4(2):161–174, April 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379571900371.

Howell:1978:GPT

[How78] Thomas D. Howell. Global properties of tensor rank. Linear Algebra and its Applications, 22(??):9–23, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900526.

Houck:1978:NSM

[HP78] David J. Houck and Michael E. Paul. Non-singular 0–1 matrices with constant row and column sums. Linear Algebra and its Applications, 22(??):263–266, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379578900769.

Houck:1979:TBE

[HP79a] David J. Houck and Michael E. Paul. On a theorem of de Bruijn and Erdős. Linear Algebra and its Applications, 23(??):157–165, February 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579900995.

Houck:1979:APN

[HP79b] David J. Houck and Arthur O. Pittenger. Analytic proofs of a network feasibility theorem and a theorem of Fulkerson. Linear Algebra and its Applications, 24(??):151–156, April 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579901551.

Hansen:1974:EGI

[HR74] Gary W. Hansen and Donald W. Robinson. On the existence of generalized inverses. Linear Algebra and its Applications, 8(2):95– 104, April 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379574900469.

Heymann:1973:NMP

[HT73] Michael Heymann and John A. Thorpe. A note on minimal polynomials. Linear Algebra and its Applications, 7(3):279–280, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379573900451.

Hu:1970:AIA

[Hu70] T. C. Hu. On the asymptotic integer algorithm. Linear Algebra and its Applications, 3(3):279–294, July 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379570900017.

Huang:1978:CPA

[Hua78] C. P. Huang. Computing powers of arbitrary Hessenberg matrices. Linear Algebra and its Applications, 21(2):123–134, August 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900368.

Haynsworth:1979:GIC

[HW79a] Emilie Haynsworth and J. R. Wall. Group inverses of certain nonnegative matrices. Linear Algebra and its Applications, 25(??):271– 288, June 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379579900235.

Hoskins:1979:FMS

[HW79b] W. D. Hoskins and D. J. Walton. A faster, more stable method for computing the p th roots of positive definite matrices. Linear Algebra and its Applications, 26(??):139–163, August 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901769.

Hoskins:1979:MSM

[HW79c] W. D. Hoskins and D. J. Walton. Methods for solving the matrix equation T B + CT = −A and its generalizations. Linear Algebra and its Applications, 23(??):217–225, February 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579901046.

Ikebe:1979:IHM

[Ike79] Yasuhiko Ikebe. On inverses of Hessenberg matrices. Linear Algebra and its Applications, 24(??):93–97, April 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579901496.

Ishaq:1979:MFA

[IM79] M. Ishaq and Russell Merris. Matrix functions afforded by entries of unitary representations. Linear Algebra and its Applications, 25(??):255–269, June 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379579900223.

Ionescu:1976:GSL

[Ion76] Tudor Ionescu. On the generators of semisimple lie algebras. Linear Algebra and its Applications, 15(3):271–292, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900331.

Joyce:1970:RIM

[JB70] G. T. Joyce and S. Barnett. Remarks on the inertia of a matrix. Linear Algebra and its Applications, 3(1):1–5, January 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379570900224.

Jensen:1973:ITR

[Jen73] Paul S. Jensen. An inclusion theorem related to inverse iteration. Linear Algebra and its Applications, 6(??):209–215, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379573900220.

Jain:1979:NTR

[JGK79] S. K. Jain, V. K. Goel, and Edward K. Kwak. Nonnegative m th roots of nonnegative 0-symmetric idempotent matrices. Linear Algebra and its Applications, 23(??):37–51, February 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579900910.

Johnson:1979:SPI

[JLR79] Charles R. Johnson, Frank Thomson Leighton, and Herbert A. Robinson. Sign patterns of inverse-positive matrices. Linear Algebra and its Applications, 24(??):75–83, April 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579901472.

Johnston:1975:BPI

[JO75] R. L. Johnston and D. D. Olesky. Best pseudo-isolated Gerschgorin disks for Eigenvalues. Linear Algebra and its Applications, 10(1):55– 67, February 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379575900944.

Johnsen:1971:EDS

[Joh71a] E. C. Johnsen. Essentially doubly stochastic matrices I. Elements of the theory over arbitrary fields. Linear Algebra and its Applications, 4(3):255–282, July 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379571900206.

Johnston:1971:GTP

[Joh71b] R. L. Johnston. Gerschgorin theorems for partitioned matrices. Linear Algebra and its Applications, 4(3):205–220, July 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379571900164.

Johnson:1973:IMW

[Joh73] Charles R. Johnson. An inequality for matrices whose symmetric part is positive definite. Linear Algebra and its Applications, 6(??): 13–18, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379573900037.

Johnsen:1974:EDS

[Joh74a] E. C. Johnsen. Essentially doubly stochastic matrices II. Canonical forms for row equivalence, equivalence, and a special case of congruence. Linear Algebra and its Applications, 8(2):121–127, April 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379574900500.

Johnson:1974:BCP

[Joh74b] James J. Johnson. Bounds for certain permanents and determinants. Linear Algebra and its Applications, 8(1):57–64, February 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379574900081.

Johnson:1974:PNP

[Joh74c] James J. Johnson. On partially non-positive matrices. Linear Algebra and its Applications, 8(2):185–187, April 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437957490055X.

Johnson:1976:LLT

[Joh76a] Charles R. Johnson. A local Lyapunov theorem and the stability of sums. Linear Algebra and its Applications, 13(1–2):37–43, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900409.

Johnson:1976:NNR

[Joh76b] Charles R. Johnson. Normality and the numerical range. Linear Algebra and its Applications, 15(1):89–94, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437957690080X.

Johnsen:1977:ESM

[Joh77a] E. C. Johnsen. Essentially stochastic matrices — factorizations into elementary and quasi-elementary matrices. Linear Algebra and its Applications, 17(1):79–93, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437957790043X.

Johnson:1977:MCG

[Joh77b] Charles R. Johnson. Multiplicativity and compatibility of generalized matrix norms. Linear Algebra and its Applications, 16(1):25–37, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900167.

Johnson:1979:PIS

[Joh79] Charles R. Johnson. Power inequalities and spectral dominance of generalized matrix norms. Linear Algebra and its Applications, 28 (??):117–130, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379579901253.

Kahane:1971:GAP

[Kah71] Joseph Kahane. Grassmann algebras for proving a theorem on Pfaffians. Linear Algebra and its Applications, 4(2):129–136, April 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379571900346.

Kaneko:1978:LCP

[Kan78a] Ikuyo Kaneko. Linear complementarity problems and characterizations of Minkowski matrices. Linear Algebra and its Applications, 20(2):111–129, ???? 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379578900459.

Kaneko:1978:RTL

[Kan78b] Ikuyo Kaneko. A reduction theorem for the linear complementarity problem with a certain patterned matrix. Linear Algebra and its Applications, 21(1):13–34, July 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379587901972.

Kapoor:1972:SMW

[Kap72] J. Kapoor. n-Square matrices which, under row permutations, leave the (n−1) th elementary symmetric function invariant. Linear Algebra and its Applications, 5(2):113–121, April 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379572900213.

Kapoor:1975:OML

[Kap75] J. Kapoor. Orthogonal matrices as linear combinations of permutation matrices. Linear Algebra and its Applications, 12(3):189–196, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379575900403.

Katz:1972:RTR

[Kat72] Irving Jack Katz. Remarks on two recent results in matrix theory. Linear Algebra and its Applications, 5(1):109–112, January 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379572900201.

Katz:1975:TPE

[Kat75] Irving J. Katz. Theorems on products of EPr matrices, II. Linear Algebra and its Applications, 10(1):37–40, February 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900920.

Kress:1974:NM

[KDW74] Rainer Kress, Hans Ludwig De Vries, and Rudolf Wegmann. On nonnormal matrices. Linear Algebra and its Applications, 8(2):109– 120, April 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379574900494.

Kellogg:1971:MSP

[Kel71] R. Bruce Kellogg. Matrices similar to a positive or essentially positive matrix. Linear Algebra and its Applications, 4(3):191–204, July 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379571900152.

Kim:1974:NIL

[Kim74] Jin Bai Kim. On the number of idempotent linear transformations. Linear Algebra and its Applications, 8(4):313–325, August 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379574900573.

Kim:1975:PLM

[Kim75] Jin Bai Kim. On r-potent linear mappings. Linear Algebra and its Applications, 10(1):69–70, February 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379575900956.

Kim:1979:EDM

[Kim79] K. H. Kim. An extension of the Dulmage–Mendelsohn theorem. Linear Algebra and its Applications, 27(??):187–197, October 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579900417.

Kostreva:1978:BPM

[Kos78] Michael M. Kostreva. Block pivot methods for solving the complementarity problem. Linear Algebra and its Applications, 21(3):207–215, September 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379578900836.

Kovarik:1974:SLB

[Kov74] Z. V. Kovarik. Spectrum localization in Banach spaces I. Linear Algebra and its Applications, 8(3):225–236, June 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379574900688.

Kovarik:1975:SLB

[Kov75] Z. V. Kovarik. Spectrum localization in Banach spaces II. Linear Algebra and its Applications, 12(3):223–229, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900440.

Kovarik:1976:CAS

[Kov76] Z. V. Kovarik. Compatibility of approximate solutions of inaccurate linear equations. Linear Algebra and its Applications, 15(3):217– 225, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379576900276.

Kovarik:1977:MCS

[Kov77] Z. V. Kovarik. Minimal compatible solutions of linear equations. Linear Algebra and its Applications, 17(2):95–106, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900301.

Kovarik:1979:MLI

[Kov79] Zdislav V. Kovarik. Manifolds of linear involutions. Linear Algebra and its Applications, 24(??):271–287, April 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579901642.

Karlin:1974:OPG

[KP74] Samuel Karlin and Allan Pinkus. Oscillation properties of generalized characteristic polynomials for totally positive and positive definite matrices. Linear Algebra and its Applications, 8(4):281–312, August 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379574900561.

Kaul:1973:TFJ

[KR73] R. N. Kaul and Oma Rani. A theorem of Fritz John in complex programming. Linear Algebra and its Applications, 7(3):217–232, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379573900414.

Kim:1978:IBM

[KR78] Ki Hang Kim and Fred W. Roush. Inverses of Boolean matrices. Linear Algebra and its Applications, 22(??):247–262, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900757.

Kim:1979:GTC

[KR79a] K. H. Kim and F. W. Roush. The group-theoretic complexity of subsemigroups of Boolean matrices. Linear Algebra and its Applications, 25(??):289–297, June 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379579900247.

Kim:1979:CER

[KR79b] Ki Hang Kim and Fred W. Roush. On a conjecture of Erdős and Rényi. Linear Algebra and its Applications, 23(??):179–189, February 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901010.

Kruskal:1977:TWA

[Kru77] Joseph B. Kruskal. Three-way arrays: rank and uniqueness of trilinear decompositions, with application to arithmetic complexity and statistics. Linear Algebra and its Applications, 18(2):95–138, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900696.

Kellogg:1978:CEN

[KS78] R. B. Kellogg and A. B. Stephens. Complex eigenvalues of a non-negative matrix with a specified graph. Linear Algebra and its Applications, 20(2):179–187, ???? 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379578900496.

Kelly:1979:MSG

[KW79] L. M. Kelly and L. T. Watson. Q-matrices and spherical geometry. Linear Algebra and its Applications, 25(??):175–189, June 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957990017X.

Kwong:1977:CUW

[Kwo77] Man Kam Kwong. Conditions under which the order of two non-negative Hermitian matrices is preserved when taking integral powers. Linear Algebra and its Applications, 18(3):223–228, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900532.

Lafon:1975:OCB

[Laf75] Jean-Claude Lafon. Optimum computation of p bilinear forms. Linear Algebra and its Applications, 10(3):225–240, June 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900713.

Laffey:1977:SQC

[Laf77] Thomas J. Laffey. Simultaneous quasidiagonalization of complex matrices. Linear Algebra and its Applications, 16(3):189–201, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900015.

Laffey:1979:NCA

[Laf79] Thomas J. Laffey. A normality criterion for an algebra of matrices. Linear Algebra and its Applications, 25(??):169–174, June 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579900168.

Lam:1975:RCS

[Lam75] Clement Wing Hong Lam. On rational circulants satisfying A^{2} = dI + λJ. Linear Algebra and its Applications, 12(2):139–150, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900622.

Langenhop:1971:LEN

[Lan71] C. E. Langenhop. The Laurent expansion for a nearly singular matrix. Linear Algebra and its Applications, 4(4):329–340, October 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379571900048.

Langenhop:1973:INS

[Lan73] Carl E. Langenhop. On the invertibility of a nearly singular matrix. Linear Algebra and its Applications, 7(4):361–365, October 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0024379573800072.

Lancaster:1977:FTLa

[Lan77a] Peter Lancaster. A fundamental theorem on lambda-matrices with applications — I. Ordinary differential equations with constant coefficients. Linear Algebra and its Applications, 18(3):189–211, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900519.

Lancaster:1977:FTLb

[Lan77b] Peter Lancaster. A fundamental theorem on lambda-matrices with applications — II. Difference equations with constant coefficients. Linear Algebra and its Applications, 18(3):213–222, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900520.

Langenhop:1977:IMP

[Lan77c] C. E. Langenhop. The inverse of a matrix polynomial. Linear Algebra and its Applications, 16(3):267–284, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437957790009X.

Lawson:1975:CMG

[Law75] Linda Marie Lawson. Computational methods for generalized inverse matrices arising from proper splittings. Linear Algebra and its Applications, 12(2):111–126, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379575900609.

Long:1970:IFC

[LC70] J. P. Long and C. G. Cullen. Intrinsic functions on complex semisimple algebras. Linear Algebra and its Applications, 3(4):429–440, October 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379570900340.

Lee:1975:BSC

[Lee75] S. L. Lee. B-splines for cardinal Hermite interpolation. Linear Algebra and its Applications, 12(3):269–280, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437957590049X.

Lehnigk:1971:HMP

[Leh71] Siegfried H. Lehnigk. On the Hermite matrix of a polynomial. Linear Algebra and its Applications, 4(4):407–419, October 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379571900127.

Levinger:1972:GCH

[Lev72a] Bernard W. Levinger. A generalization of the Camion–Hoffman theorem. Linear Algebra and its Applications, 5(3):227–232, July 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379572900043.

Levow:1972:PMC

[Lev72b] Roy B. Levow. A problem of Mirsky concerning nonsingular doubly stochastic matrices. Linear Algebra and its Applications, 5(3):197– 206, July 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/002437957290002X.

Lax:1978:SS

[LL78] Anneli Lax and Peter D. Lax. On sums of squares. Linear Algebra and its Applications, 20(1):71–75, March 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379578900319.

Laffey:1977:NSM

[LM77] Thomas J. Laffey and Donald L. McQuillan. A note on subfields of matrix rings. Linear Algebra and its Applications, 16(1):1–3, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900131.

Lau:1978:FNM

[LM78] Cony M. Lau and Thomas L. Markham. Factorization of nonnegative matrices — II. Linear Algebra and its Applications, 20(1):51– 56, March 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379578900289.

Lorenz:1979:TMT

[LM79] Jens Lorenz and Wolfgang Mackens. Toeplitz matrices with totally nonnegative inverses. Linear Algebra and its Applications, 24(??):133– 141, April 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379579901538.

Lovass–Nagy:1970:FPM

[LNP70] Victor Lovass-Nagy and David L. Powers. On functions of partitioned matrices with nondiagonalizable submatrices. Linear Algebra and its Applications, 3(2):257–262, April 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379570900182.

Lovass-Nagy:1971:CRI

[LNP71] V. Lovass-Nagy and D. L. Powers. On the commuting reciprocal inverse of some partitioned matrices. Linear Algebra and its Applications, 4(2):183–190, April 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379571900395.

Lovass-Nagy:1973:MAM

[LNPU73] V. Lovass-Nagy, D. L. Powers, and F. D. Ullman. A modified ADI method for computing the “best least-squares solution” of an incompatible system (A × I + I × B)x = g. Linear Algebra and its Applications, 7(2):179–185, April 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379573900530.

Lim:1978:LTS

[LO78] M. H. Lim and Hock Ong. Linear transformations on symmetric matrices that preserve the permanent. Linear Algebra and its Applications, 21(2):143–151, August 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379578900381.

Loewy:1971:TAL

[Loe71] Raphael Loewy. On a theorem about the location of eigenvalues of matrices. Linear Algebra and its Applications, 4(3):233–242, July 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379571900188.

Loewy:1976:RRL

[Loe76] Raphael Loewy. On ranges of real Lyapunov transformations. Linear Algebra and its Applications, 13(1–2):79–89, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900458.

London:1971:SNV

[Lon71] David London. Some notes on the van der Waerden conjecture. Linear Algebra and its Applications, 4(2):155–160, April 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957190036X.

Longstaff:1979:OAL

[Lon79] W. E. Longstaff. On the operation Alg Lat in finite dimensions. Linear Algebra and its Applications, 27(??):27–29, October 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579900296.

Lipow:1973:CIS

[LS73] Peter R. Lipow and I. J. Schoenberg. Cardinal interpolation and spline functions. III. Cardinal Hermite interpolation. Linear Algebra and its Applications, 6(??):273–304, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379573900293.

Loewy:1975:IC

[LS75] Raphael Loewy and Hans Schneider. Indecomposable cones. Linear Algebra and its Applications, 11(3):235–245, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900233.

Loughlin:1976:MPM

[LS76] Timothy A. Loughlin and Paul Slepian. Minimal paramount matrices. Linear Algebra and its Applications, 15(3):197–204, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900252.

Long:1975:FFF

[LV75] Andrew F. Long, Jr. and Theresa P. Vaughan. Factorization of Q(h(T)(x)) over a finite field where Q(x) is irreducible and h(T)(x) is linear II. Linear Algebra and its Applications, 11(1):53–72, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575901172.

Long:1976:FFF

[LV76] Andrew F. Long and Theresa P. Vaughan. Factorization of Q(h(T)(x)) over a finite field where Q(x) is irreducible and h(T)(x) is linear I. Linear Algebra and its Applications, 13(3):207–221, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900975.

Lyle:1978:LAP

[Lyl78] Bill Lyle. A linear-algebra problem from algebraic coding theory. Linear Algebra and its Applications, 22(??):223–233, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900733.

Marcus:1977:NRE

[MA77] Marvin Marcus and Patricia Andresen. The numerical radius of exterior powers. Linear Algebra and its Applications, 16(2):131–151, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957790026X.

Mees:1979:DCF

[MA79] A. I. Mees and D. P. Atherton. Domains containing the field of values of a matrix. Linear Algebra and its Applications, 26(??):289–296, August 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379579901836.

Majindar:1979:LCH

[Maj79] K. N. Majindar. Linear combinations of Hermitian and real symmetric matrices. Linear Algebra and its Applications, 25(??):95–105, June 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579900090.

Mangasarian:1974:MSM

[Man74] O. L. Mangasarian. Monotone Splitting of matrices. Linear Algebra and its Applications, 8(1):43–55, February 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437957490007X.

Mangasarian:1979:USL

[Man79] O. L. Mangasarian. Uniqueness of solution in linear programming. Linear Algebra and its Applications, 25(??):151–162, June 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579900144.

Markham:1970:OM

[Mar70a] Thomas L. Markham. On oscillatory matrices. Linear Algebra and its Applications, 3(2):143–156, April 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379570900108.

Markham:1970:STN

[Mar70b] Thomas L. Markham. A semigroup of totally nonnegative matrices. Linear Algebra and its Applications, 3(2):157–164, April 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957090011X.

Marshall:1973:NIC

[Mar73] Albert W. Marshall. Norms and inequalities for condition numbers, III. Linear Algebra and its Applications, 7(4):291–300, October 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0024379573800023.

Marcus:1975:RER

[Mar75] Marvin Marcus. Rearrangement and extremal results for Hermitian matrices. Linear Algebra and its Applications, 11(2):95–104, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957590052X.

Marcus:1979:VCS

[Mar79a] Marvin Marcus. Variations on the Cauchy–Schwarz inequality. Linear Algebra and its Applications, 27(??):81–91, October 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579900338.

Markham:1979:TPM

[Mar79b] Thomas L. Markham. Two properties of M-matrices. Linear Algebra and its Applications, 28(??):131–134, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901265.

MarquesdeSa:1979:ICM

[Mar79c] E. Marques de Sá. Imbedding conditions for λ-matrices. Linear Algebra and its Applications, 24(??):33–50, April 1979. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0024379579901459.

MarquesdeSa:1979:EMC

[Mar79d] E. Marques de Sá. On the existence of matrices with constrained elements, rows and columns. Linear Algebra and its Applications, 24(??):263–270, April 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379579901630.

Mattheij:1975:AES

[Mat75] R. M. M. Mattheij. Accurate estimates of solutions of second order recursions. Linear Algebra and its Applications, 12(1):29–54, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575901251.

Maurer:1975:MRM

[Mau75] Stephen B. Maurer. A maximum-rank minimum-term-rank theorem for matroids. Linear Algebra and its Applications, 10(2):129– 137, April 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/002437957590004X.

Maxwell:1971:DC

[Max71] George Maxwell. On double commutators. Linear Algebra and its Applications, 4(3):283–284, July 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379571900218.

Maroulas:1978:GCD

[MB78] John Maroulas and Stephen Barnett. Greatest common divisor of generalized polynomials and polynomial matrices. Linear Algebra and its Applications, 22(??):195–210, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002437957890071X.

McDonald:1978:SMA

[McD78] B. R. McDonald. Similarity of matrices over Artinian principal ideal rings. Linear Algebra and its Applications, 21(2):153–162, August 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900393.

Meek:1976:NCM

[Mee76] D. S. Meek. A new class of matrices with positive inverses. Linear Algebra and its Applications, 15(3):253–260, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900318.

Merris:1971:GAT

[Mer71] Russel Merris. A generalization of the associated transformation. Linear Algebra and its Applications, 4(4):393–406, October 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379571900115.

Merris:1975:TPI

[Mer75] Russell Merris. Two problems involving Schur functions. Linear Algebra and its Applications, 10(2):155–162, April 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900075.

Merris:1977:NDS

[Mer77] Russell Merris. Nonzero decomposable symmetrized tensors. Linear Algebra and its Applications, 17(3):287–292, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900672.

Merris:1978:ASF

[Mer78] Russell Merris. Applications of spherical functions in multilinear algebra. Linear Algebra and its Applications, 20(1):77–86, March 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900320.

Metelmann:1973:KNT

[Met73] Kurt Metelmann. Ein Kriterium für den Nachweis der Totalnichtnegativität von Bandmatrizen. (German) [A criterion for the detection of total non-negativity of band matrices]. Linear Algebra and its Applications, 7(2):163–171, April 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379573900505.

Meyer:1970:SRE

[Mey70] Carl D. Meyer, Jr. Some remarks on EPγ matrices, and generalized inverses. Linear Algebra and its Applications, 3(2):275–278, April 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379570900200.

Meyer:1971:RIP

[Mey71] Carl D. Meyer, Jr. Representations for (1)- and (1,2)-inverses for partitioned matrices. Linear Algebra and its Applications, 4(3):221– 232, July 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379571900176.

Meyer:1978:MRS

[Mey78a] Hans-Bernhard Meyer. Matrix Riccati solutions. Linear Algebra and its Applications, 20(2):131–146, ???? 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379578900460.

Meyer:1978:AEM

[Mey78b] Carl D. Meyer, Jr. An alternative expression for the mean first passage matrix. Linear Algebra and its Applications, 22(??):41–47, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900551.

Marcus:1978:NBP

[MF78] Marvin Marcus and Ivan Filippenko. Nondifferentiable boundary points of the higher numerical range. Linear Algebra and its Applications, 21(3):217–232, September 1978. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379578900848.

Marcus:1970:GUG

[MG70] Marvin Marcus and William Robert Gordon. A generalization of the unitary group. Linear Algebra and its Applications, 3(2):225– 247, April 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379570900169.

Mihalyffy:1971:ARG

[Mih71] L. Mihályffy. An alternative representation of the generalized inverse of partitioned matrices. Linear Algebra and its Applications, 4(1):95– 100, January 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379571900310.

Minc:1972:NDM

[Min72] Henryk Minc. Nearly decomposable matrices. Linear Algebra and its Applications, 5(2):181–187, April 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379572900274.

Minc:1979:CRF

[Min79] Henryk Minc. On a conjecture of R. F. Scott (1881). Linear Algebra and its Applications, 28(??):141–153, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901289.

Mlynarski:1977:NMM

[Mly77] Max Mlynarski. A note on matricial minimality. Linear Algebra and its Applications, 17(2):133–138, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379577900337.

Marcus:1972:AES

[MM72a] Marvin Marcus and Russell Merris. Antiderivations on the exterior and symmetric algebras. Linear Algebra and its Applications, 5(1):13– 18, January 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379572900146.

Marcus:1972:ISF

[MM72b] Marvin Marcus and Henryk Minc. An inequality for Schur functions. Linear Algebra and its Applications, 5(1):19–28, January 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379572900158.

Marcus:1970:SRU

[MN70] Marvin Marcus and Morris Newman. Some results on unitary matrix groups. Linear Algebra and its Applications, 3(2):173–178, April 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379570900133.

McCormick:1977:SIM

[MN77] S. F. McCormick and T. Noe. Simultaneous iteration for the matrix eigenvalue problem. Linear Algebra and its Applications, 16(1):43– 56, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379577900180.

Masser:1979:SRS

[MN79] D. W. Masser and M. Neumann. On the square roots of strictly quasiaccretive complex matrices. Linear Algebra and its Applications, 28(??):135–140, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379579901277.

Mohan:1976:SMC

[Moh76] S. R. Mohan. On the simplex method and a class of linear complementary problems. Linear Algebra and its Applications, 14(1):1–9, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900574.

Moore:1974:VPF

[Moo74] Michael H. Moore. Vector packing in finite dimensional vector spaces. Linear Algebra and its Applications, 8(3):213–224, June 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379574900676.

Morel:1976:APP

[Mor76] Paul Morel. Des algorithmes pour le problème inverse des valeurs propres. (French) [Algorithms for the inverse problem of eigenvalues]. Linear Algebra and its Applications, 13(3):251–273, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576901002.

Mizori-Oblak:1978:EMV

[MOV78] Pavla Mizori-Oblak and Ivan Vidav. On entire matrix-valued functions. Linear Algebra and its Applications, 22(??):25–31, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900538.

Moylan:1977:MPP

[Moy77] Peter J. Moylan. Matrices with positive principal minors. Linear Algebra and its Applications, 17(1):53–58, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379577900404.

Mitra:1979:SOG

[MP79] Sujit Kumar Mitra and Madan L. Puri. Shorted operators and generalized inverses of matrices. Linear Algebra and its Applications, 25(??):45–56, June 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379579900053.

More:1973:FRC

[MR73] J. Moré and W. Rheinboldt. On P - and S-functions and related classes of n-dimensional nonlinear mappings. Linear Algebra and its Applications, 6(??):45–68, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379573900062.

Marcus:1975:NHS

[MR75] Marvin Marcus and Herbert Robinson. A note on the Hodge star operator. Linear Algebra and its Applications, 10(1):85–87, February 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900981.

Marcus:1976:EPE

[MR76] Marvin Marcus and Herbert Robinson. On exterior powers of endomorphisms. Linear Algebra and its Applications, 14(3):219–225, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900689.

Meyer:1978:SMI

[MS78] C. D. Meyer, Jr. and M. W. Stadelmaier. Singular M-matrices and inverse positivity. Linear Algebra and its Applications, 22(??):139–156, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379578900654.

McMath:1977:PMV

[MSH77] John S. McMath, S. E. Sins, and C. R. Hallum. On the properties of a matrix valued derivative utilizing the Moore–Penrose inverse. Linear Algebra and its Applications, 18(3):281–291, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957790057X.

Mitra:1975:HAM

[MT75] Sujit Kumar Mitra and George E. Trapp. On hybrid addition of matrices. Linear Algebra and its Applications, 10(1):19–35, February 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900919.

Marchi:1979:ASM

[MT79] Ezio Marchi and Pablo Tarazaga. About (k, n) stochastic matrices. Linear Algebra and its Applications, 26(??):15–30, August 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901691.

Muller:1973:SVR

[Mül73] Karl Hans Müller. Über den spektralradius verallgemeinerter relaxationsverfahren. (German) [On the spectral-radius generalized relaxation method]. Linear Algebra and its Applications, 6(??):263–268, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/002437957390027X.

Mullen:1979:ECM

[Mul79] Gary L. Mullen. Equivalence classes of matrices over finite fields. Linear Algebra and its Applications, 27(??):61–68, October 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579900314.

Murty:1972:NSC

[Mur72] Katta G. Murty. On the number of solutions to the complementarity problem and spanning properties of complementary cones. Linear Algebra and its Applications, 5(1):65–108, January 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379572900195.

Micchelli:1979:FWP

[MW79] Charles A. Micchelli and R. A. Willoughby. On functions which preserve the class of Stieltjes matrices. Linear Algebra and its Applications, 23(??):141–156, February 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579900983.

Natterer:1973:GSM

[Nat73] Frank Natterer. A generalized spline method for singular boundary value problems of ordinary differential equations. Linear Algebra and its Applications, 7(3):189–216, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379573900402.

Nazareth:1975:CCJ

[Naz75] Larry Nazareth. On the convergence of the cyclic Jacobi method. Linear Algebra and its Applications, 12(2):151–164, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900634.

Nazareth:1977:UAU

[Naz77] Larry Nazareth. Unified approach to unconstrained minimization via basic matrix factorizations. Linear Algebra and its Applications, 17(3):197–232, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379577900623.

Neumann:1976:SSR

[Neu76] M. Neumann. Subproper splitting for rectangular matrices. Linear Algebra and its Applications, 14(1):41–51, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900628.

Neumann:1978:SAP

[Neu78] M. Neumann. Some applications of partial orderings to iterative methods for rectangular linear systems. Linear Algebra and its Applications, 19(2):95–116, ???? 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379578900095.

Neumann:1979:NGS

[Neu79] M. Neumann. A note on generalizations of strict diagonal dominance for real matrices. Linear Algebra and its Applications, 26(??):3–14, August 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/002437957990168X.

Ng:1976:ECC

[Ng76] Dina Ng. An effective criterion for congruence of real symmetric matrix pairs. Linear Algebra and its Applications, 13(1–2):11–18, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900380.

Nicholson:1975:MPE

[Nic75] Victor A. Nicholson. Matrices with permanent equal to one. Linear Algebra and its Applications, 12(2):185–188, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900671.

Nicholson:1979:EBP

[Nic79] D. W. Nicholson. Eigenvalue bounds for AB+BA, with A, B positive definite matrices. Linear Algebra and its Applications, 24(??):173– 184, April 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379579901575.

Neumann:1979:GIP

[NP79] M. Neumann and R. J. Plemmons. Generalized inverse-positivity and splittings of M-matrices. Linear Algebra and its Applications, 23(??):21–35, February 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379579900909.

Nguyen:1979:SPU

[NS79] V. Hien Nguyen and J.-J. Strodiot. Some properties of a uniformly linearly independent sequence of subspaces. Linear Algebra and its Applications, 25(??):115–127, June 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579900119.

Nwokah:1978:EIM

[Nwo78] Osita D. I. Nwokah. Estimates for the inverse of a matrix and bounds for eigenvalues. Linear Algebra and its Applications, 22(??):283–291, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379578900782.

OBrien:1975:EME

[O’B75] Richard C. O’Brien. On extreme matrices and extreme vectors of cones in R^{n}. Linear Algebra and its Applications, 12(1):77–79, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575901287.

Okoh:1979:DSD

[Oko79] Frank Okoh. Direct sums and direct products of canonical pencils of matrices. Linear Algebra and its Applications, 25(??):1–26, June 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579900028.

Olkin:1979:MEL

[Olk79] Ingram Olkin. Matrix extensions of Liouville–Dirichlet-type integrals. Linear Algebra and its Applications, 28(??):155–160, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901290.

Ong:1976:LTM

[Ong76] Hock Ong. Linear transformations on matrices: the invariance of generalized permutation matrices — III. Linear Algebra and its Applications, 15(2):119–151, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379576900124.

Ortega:1979:EOR

[OP79] James M. Ortega and Robert J. Plemmons. Extensions of the Ostrowski–Reich theorem for SOR iterations. Linear Algebra and its Applications, 28(??):177–191, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579901319.

OReilly:1977:MAD

[O’R77] Michael F. O’Reilly. Matrix algebras with direct product. Linear Algebra and its Applications, 16(1):19–24, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900155.

Olkin:1972:JMT

[OS72] Ingram Olkin and Allan R. Sampson. Jacobians of matrix transformations and induced functional equations. Linear Algebra and its Applications, 5(3):257–276, July 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379572900079.

Ostrowski:1971:NPH

[Ost71] Alexander Ostrowski. A new proof of Haynsworth’s quotient formula for Schur complements. Linear Algebra and its Applications, 4(4):389– 392, October 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379571900103.

Ostrowski:1974:SRN

[Ost74] A. M. Ostrowski. On subdominant roots of nonnegative matrices. Linear Algebra and its Applications, 8(2):179–184, April 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379574900548.

Padberg:1976:AIP

[Pad76] Manfred W. Padberg. Almost integral polyhedra related to certain combinatorial optimization problems. Linear Algebra and its Applications, 15(1):69–88, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379576900793.

Paige:1974:EPH

[Pai74] C. C. Paige. Eigenvalues of perturbed Hermitian matrices. Linear Algebra and its Applications, 8(1):1–10, February 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379574900020.

Pall:1972:PMM

[Pal72] Gordon Pall. Partitioning by means of maximal isotropic subspaces. Linear Algebra and its Applications, 5(2):173–180, April 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379572900262.

Pang:1978:COL

[Pan78] Jong-Shi Shi Pang. On cone orderings and the linear complementarity problem. Linear Algebra and its Applications, 22(??):267–281, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379578900770.

Pang:1979:HMP

[Pan79a] Jong-Shi Pang. Hidden Z-matrices with positive principal minors. Linear Algebra and its Applications, 23(??):201–215, February 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901034.

Pang:1979:NCR

[Pan79b] Jong-Shi Shi Pang. A new characterization of real H-matrices with positive diagonals. Linear Algebra and its Applications, 25(??):163– 167, June 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379579900156.

Pao:1973:FRL

[Pao73a] C. V. Pao. A further remark on the logarithmic derivatives of a square matrix. Linear Algebra and its Applications, 7(3):275–278, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/002437957390044X.

Pao:1973:LDS

[Pao73b] C. V. Pao. Logarithmic derivatives of a square matrix. Linear Algebra and its Applications, 6(??):159–164, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379573900153.

Parsons:1970:CAC

[Par70] T. D. Parsons. A combinatorial approach to convex quadratic programming. Linear Algebra and its Applications, 3(3):359–378, July 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379570900066.

Parlett:1973:NHM

[Par73] Beresford Parlett. Normal Hessenberg and moment matrices. Linear Algebra and its Applications, 6(??):37–43, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379573900050.

Parameswaran:1975:RBC

[Par75] M. R. Parameswaran. Relations between certain matrix products similar to commutativity. Linear Algebra and its Applications, 10(3):219– 224, June 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379575900701.

Parlett:1976:RAE

[Par76] B. N. Parlett. A recurrence among the elements of functions of triangular matrices. Linear Algebra and its Applications, 14(2):117– 121, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379576900185.

Partington:1979:HOA

[Par79] J. R. Partington. Hermitian operators for absolute norms and absolute direct sums. Linear Algebra and its Applications, 23(??):275–280, February 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379579901095.

Pate:1976:LBN

[Pat76] Thomas H. Pate. Lower bounds for the norm of the symmetric product. Linear Algebra and its Applications, 14(3):285–292, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900732.

Paz:1975:WFP

[Paz75] A. Paz. Word functions of pseudo-Markov chains. Linear Algebra and its Applications, 10(1):1–5, February 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379575900890.

Poole:1979:MM

[PB79] G. D. Poole and G. P. Barker. MP matrices. Linear Algebra and its Applications, 26(??):165–174, August 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579901770.

Pye:1973:PCM

[PBA73] W. C. Pye, T. L. Boullion, and T. A. Atchison. The pseudoinverse of a centrosymmetric matrix. Linear Algebra and its Applications, 6(??):201–204, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379573900207.

Pinkus:1979:MW

[Pin79] Allan Pinkus. Matrices and n-Widths. Linear Algebra and its Applications, 27(??):245–278, October 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579900466.

Pye:1975:MPL

[PJ75] W. C. Pye and Melvyn W. Jeter. A minimal permanent-like function. Linear Algebra and its Applications, 12(2):171–178, ????

1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900658.

Plemmons:1972:GNM

[Ple72] R. J. Plemmons. Graphs and nonnegative matrices. Linear Algebra and its Applications, 5(3):283–292, July 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379572900092.

Plemmons:1976:MLS

[Ple76] R. J. Plemmons. M-matrices leading to semiconvergent splittings. Linear Algebra and its Applications, 15(3):243–252, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900306.

Plemmons:1977:MCN

[Ple77] R. J. Plemmons. M-matrix characterizations. I. Nonsingular Mmatrices. Linear Algebra and its Applications, 18(2):175–188, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900738.

Porter:1979:RS

[PM79] A. Duane Porter and Nick Mousouris. Ranked solutions of AXC = B and AX = B. Linear Algebra and its Applications, 24(??):217– 224, April 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379579901599.

Porter:1975:SPH

[Por75] A. Duane Porter. Some partitions of a Hermitian matrix. Linear Algebra and its Applications, 12(3):231–239, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900452.

Porter:1976:SME

[Por76] A. Duane Porter. Solvability of the matric equation AX = B. Linear Algebra and its Applications, 13(3):177–184, ???? 1976.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957690094X.

Powers:1976:BGT

[Pow76] David L. Powers. A block Geršgorin theorem. Linear Algebra and its Applications, 13(1–2):45–52, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379576900410.

Puri:1977:MMW

[PR77] Prem S. Puri and V. Ramaswami. Maximum and minimum weighted diagonal sums of certain non-negative matrices. Linear Algebra and its Applications, 16(3):247–266, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379577900088.

Prather:1976:EFM

[Pra76] Ronald E. Prather. Epimorphisms of free monoids. Linear Algebra and its Applications, 13(3):201–205, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379576900963.

Patel:1979:TII

[PT79a] Rajnikant Patel and Mitsuhiko Toda. Trace inequalities involving Hermitian matrices. Linear Algebra and its Applications, 23(??):13– 20, February 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379579900892.

Phadke:1979:GIO

[PT79b] S. V. Phadke and N. K. Thakare. Generalized inverses and operator equations. Linear Algebra and its Applications, 23(??):191–199, February 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379579901022.

Ptak:1976:NMI

[Ptá76] Vlastimil Pták. Nondiscrete mathematical induction and iterative existence proofs. Linear Algebra and its Applications, 13(3):223– 238, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379576900987.

Ptak:1979:MPM

[Ptá79] Vlastimil Pták. A maximum problem for matrices. Linear Algebra and its Applications, 28(??):193–204, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579901320.

Pullman:1971:GAT

[Pul71] N. J. Pullman. A geometric approach to the theory of nonnegative matrices. Linear Algebra and its Applications, 4(4):297–312, October 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379571900012.

Putcha:1979:QSA

[Put79] Mohan S. Putcha. Quadratic semigroups on affine spaces. Linear Algebra and its Applications, 26(??):107–121, August 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901745.

Quackenbush:1975:NVS

[Qua75] Robert W. Quackenbush. Near vector spaces over GF(q) and (v, q + 1, 1)-BIBD’s. Linear Algebra and its Applications, 10(3):259– 266, June 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379575900737.

Radjavi:1975:DMS

[Rad75] Heydar Radjavi. Decomposition of matrices into simple involutions. Linear Algebra and its Applications, 12(3):247–255, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900476.

Raghavan:1978:CMG

[Rag78] T. E. S. Raghavan. Completely mixed games and M-matrices. Linear Algebra and its Applications, 21(1):35–45, July 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379587901984.

Robers:1970:SMI

[RBI70] Philip D. Robers and Adi Ben-Israel. A suboptimization method for interval linear programming: a new method for linear programming. Linear Algebra and its Applications, 3(3):383–405, July 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957090008X.

Rebman:1974:TUT

[Reb74] Kenneth R. Rebman. Total unimodularity and the transportation problem: a generalization. Linear Algebra and its Applications, 8 (1):11–24, February 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379574900032.

Rehm:1977:OSM

[Reh77] Hans P. Rehm. On Ochoa’s special matrices in matrix classes. Linear Algebra and its Applications, 17(2):181–188, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900362.

Robinson:1971:SEH

[Rob71] Donald W. Robinson. Separable endomorphisms and higher-order commutators. Linear Algebra and its Applications, 4(3):285–296, July 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957190022X.

Robinson:1973:BES

[Rob73] Stephen M. Robinson. Bounds for error in the solution set of a perturbed linear program. Linear Algebra and its Applications, 6(??): 69–81, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379573900074.

Robbin:1975:LAI

[Rob75] J. W. Robbin. Lie algebras of infinitesimal norm isometries. Linear Algebra and its Applications, 10(2):95–102, April 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900014.

Robert:1976:CNV

[Rob76] F. Robert. Contraction en norme vectorielle: Convergence d’itérations chaotiques pour des équations non linéaires de point fixe à plusieurs variables. (French) [Contraction in vector norm: Convergence of chaotic iterations for nonlinear equations of fixed point in several variables]. Linear Algebra and its Applications, 13(1–2): 19–35, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379576900392.

Robert:1978:TPF

[Rob78a] François Robert. Théorèmes de Perron–Frobenius et Stein–Rosenberg booléens. (French) [Perron–Frobenius and Stein–Rosenberg Boolean theorems]. Linear Algebra and its Applications, 19(3):237–250, ???? 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900137.

Robinson:1978:MTM

[Rob78b] Herbert A. Robinson. Multilinear transformation on matrices. Linear Algebra and its Applications, 20(3):205–218, June 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900198.

Roitman:1978:PCM

[Roi78] Moshe Roitman. A problem on conjugacy of matrices. Linear Algebra and its Applications, 19(1):87–89, ???? 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379578900071.

Roscoe:1976:AMT

[Ros76a] D. F. Roscoe. The application of matrix theorems to the construction of stable difference representations of the equations of fluid dynamics. Linear Algebra and its Applications, 14(2):123–130, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900197.

Rose:1976:NCD

[Ros76b] Nicholas J. Rose. A note on computing the Drazin inverse. Linear Algebra and its Applications, 15(2):95–98, ???? 1976.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900100.

Rothblum:1975:AEN

[Rot75] Uriel G. Rothblum. Algebraic eigenspaces of nonnegative matrices. Linear Algebra and its Applications, 12(3):281–292, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900506.

Rothblum:1979:ICM

[Rot79a] Uriel G. Rothblum. An index classification of M-matrices. Linear Algebra and its Applications, 23(??):1–12, February 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579900880.

Rothblum:1979:RCN

[Rot79b] Uriel G. Rothblum. A rank characterization of the number of final classes of a nonnegative matrix. Linear Algebra and its Applications, 23(??):65–68, February 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379579900934.

Robert:1973:CFN

[RR73] F. Robert and M. Rascle. Contraction faible en norme vectorielle — théorie de Perron–Frobenius pour le cas de blocs. (French) [Weak contraction in standard vector norm — Perron–Frobenius theory for the case of blocks]. Linear Algebra and its Applications, 6(??):305– 335, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/002437957390030X.

Rao:1975:GIB

[RR75] P. S. S. N. V. Prasada Rao and K. P. S. Bhaskara Rao. On generalized inverses of Boolean Matrices. Linear Algebra and its Applications, 11(2):135–153, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379575900543.

Robinson:1971:GPM

[RS71] Norman F. Robinson and M. F. Smiley. A generalization of the property L of Motzkin and Taussky. Linear Algebra and its Applications, 4(4):323–328, October 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379571900036.

Rheinboldt:1974:CTM

[RS74] Werner C. Rheinboldt and Roger A. Shepherd. On a characterization of tridiagonal matrices by M. Fiedler. Linear Algebra and its Applications, 8(1):87–90, February 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379574900111.

Ruhe:1975:CES

[Ruh75] Axel Ruhe. On the closeness of eigenvalues and singular values for almost normal matrices. Linear Algebra and its Applications, 11(1): 87–93, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379575901196.

Rackusin:1977:LMM

[RW77] Jeffrey Rackusin and William Watkins. Linear maps on matrices: the invariance of symmetric polynomials. Linear Algebra and its Applications, 17(3):269–276, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379577900647.

Ryavec:1979:IZD

[Rya79] C. Ryavec. Inequalities for the zeros of Dirichlet L-functions. Linear Algebra and its Applications, 28(??):205–211, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901332.

Ryser:1970:GME

[Rys70] H. J. Ryser. A generalization of the matrix equation A^{2} = J. Linear Algebra and its Applications, 3(4):451–460, October 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379570900364.

Saigal:1972:MM

[Sai72] Romesh Saigal. On the modularity of a matrix. Linear Algebra and its Applications, 5(1):39–48, January 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379572900171.

Sankoff:1972:CFG

[San72] David Sankoff. Context-free grammars and nonnegative matrices. Linear Algebra and its Applications, 5(3):277–281, July 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379572900080.

Saunders:1977:CCN

[Sau77] B. David Saunders. A condition for the convexity of the normnumerical range of a matrix. Linear Algebra and its Applications, 16(2):167–175, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379577900283.

Schneider:1970:BN

[Sch70] Hans Schneider. Bound norms. Linear Algebra and its Applications, 3(1):11–21, January 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379570900248.

Schneider:1977:CIF

[Sch77] Hans Schneider. The concepts of irreducibility and full indecomposability of a matrix in the works of Frobenius, König and Markov. Linear Algebra and its Applications, 18(2):139–162, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900702.

Scheffold:1978:ASE

[Sch78] Egon Scheffold. Eine Abschätzung für die subdominanten eigenwerte nichtnegativer matrizen. (German) [An estimate for the subdominant eigenvalues of nonnegative matrices]. Linear Algebra and its Applications, 19(1):91–93, ???? 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379578900083.

Schonhage:1979:APH

[Sch79] Arnold Schönhage. Arbitrary perturbations of Hermitian matrices. Linear Algebra and its Applications, 24(??):143–149, April 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957990154X.

Sharma:1977:LCC

[SD77] Bhu Dev Sharma and Bal Kishan Dass. On linear codes correcting random and low-density burst errors. Linear Algebra and its Applications, 16(1):5–18, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379577900143.

Searle:1979:ICM

[Sea79] S. R. Searle. On inverting circulant matrices. Linear Algebra and its Applications, 25(??):77–89, June 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579900077.

Seidman:1974:ICM

[Sei74] Thomas I. Seidman. An inequality for complex matrices. Linear Algebra and its Applications, 8(4):327–330, August 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379574900585.

Sengupta:1975:CEI

[Sen75] Sailes Kumar Sengupta. Comparison of eigenvectors of irreducible stochastic matrices. Linear Algebra and its Applications, 12(2):101– 110, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379575900592.

Sharpe:1977:GTL

[Sha77] D. W. Sharpe. Grade and the theory of linear equations. Linear Algebra and its Applications, 18(1):25–32, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379577900763.

Shapiro:1979:SBT

[Sha79] H. Shapiro. Simultaneous block triangularization and block diagonalization of sets of matrices. Linear Algebra and its Applications, 25(??):129–137, June 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379579900120.

Sierksma:1979:NMO

[Sie79] Gerard Sierksma. Nonnegative matrices: the open Leontief model. Linear Algebra and its Applications, 26(??):175–201, August 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901782.

Silber:1970:LDT

[Sil70] Robert Silber. A linear dependence theorem with application to the calculus of variations. Linear Algebra and its Applications, 3(4):441– 450, October 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379570900352.

Sinkhorn:1977:DSM

[Sin77] Richard Sinkhorn. Doubly stochastic matrices which have certain diagonals with constant sums. Linear Algebra and its Applications, 16(1):79–82, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379577900210.

Sinkhorn:1979:RBS

[Sin79] Richard Sinkhorn. A relationship between stochastic matrices and eigenvalues of products. Linear Algebra and its Applications, 23 (??):53–63, February 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379579900922.

Sizer:1977:TSM

[Siz77] Walter S. Sizer. Triangularizing semigroups of matrices over a skew field. Linear Algebra and its Applications, 16(2):177–187, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900295.

Smith:1971:ACH

[Smi71] Russell A. Smith. Area of convex hull of matrix eigenvalues. Linear Algebra and its Applications, 4(1):77–78, January 1971.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379571900292.

Smith:1974:CMR

[Smi74] Kirby C. Smith. The centralizer of a matrix with real quaternion elements. Linear Algebra and its Applications, 8(4):331–335, August 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379574900597.

Smith:1976:CSL

[Smi76] Kirby C. Smith. The centralizer of a semi-linear transformation. Linear Algebra and its Applications, 14(1):21–27, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900598.

Smith:1977:MPI

[Smi77] Ronald L. Smith. Moore–Penrose inverses of block circulant and block k-circulant matrices. Linear Algebra and its Applications, 16(3):237– 245, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379577900076.

Smith:1978:MPI

[Smi78] Ronald L. Smith. The Moore–Penrose inverse of a retrocirculant. Linear Algebra and its Applications, 22(??):1–8, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900514.

Smith:1979:FSS

[Smi79] Ronald L. Smith. Factorization of symmetric singular M-matrices. Linear Algebra and its Applications, 23(??):253–259, February 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901071.

Spain:1978:SOE

[Spa78] P. G. Spain. Solving an operator equation by iteration. Linear Algebra and its Applications, 20(1):69–70, March 1978. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379578900307.

Schoenberg:1973:CIS

[SS73] I. J. Schoenberg and A. Sharma. Cardinal interpolation and spline functions V. The B-splines for cardinal Hermite interpolation. Linear Algebra and its Applications, 7(1):1–42, January 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379573900347.

Schmidt:1976:BMI

[SS76] Gunther Schmidt and Thomas Ströhlein. A Boolean matrix iteration in timetable construction. Linear Algebra and its Applications, 15(1): 27–51, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/002437957690077X.

Shinozaki:1979:FRR

[SS79] Nobuo Shinozaki and Masaaki Sibuya. Further results on the reverseorder law. Linear Algebra and its Applications, 27(??):9–16, October 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579900272.

Standish:1975:TTS

[Sta75] C. J. Standish. Truncated Taylor series approximation to the state transition matrix of a continuous parameter finite Markov chain. Linear Algebra and its Applications, 12(2):179–183, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957590066X.

Stafney:1978:FMT

[Sta78] James D. Stafney. Functions of a matrix and their norms. Linear Algebra and its Applications, 20(1):87–94, March 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900332.

Stein:1977:MLD

[Ste77] S. K. Stein. Modified linear dependence and the capacity of a cyclic graph. Linear Algebra and its Applications, 17(3):191–195, ????

1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900611.

Stewart:1979:NPS

[Ste79a] G. W. Stewart. A note on the perturbation of singular values. Linear Algebra and its Applications, 28(??):213–216, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901344.

Stewart:1979:PBD

[Ste79b] G. W. Stewart. Perturbation bounds for the definite generalized eigenvalue problem. Linear Algebra and its Applications, 23(??):69–85, February 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379579900946.

Stewart:1979:PH

[Ste79c] G. W. Stewart. Preface: A. S. Householder. Linear Algebra and its Applications, 28(??):1–3 + 1, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579901125.

Stoer:1979:TEC

[Sto79] J. Stoer. Two examples on the convergence of certain rank-2 minimization methods for quadratic functionals in Hilbert space. Linear Algebra and its Applications, 28(??):217–222, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901356.

Styan:1973:HPM

[Sty73] George P. H. Styan. Hadamard products and multivariate statistical analysis. Linear Algebra and its Applications, 6(??):217–240, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379573900232.

Shivakumar:1973:LEI

[SW73] P. N. Shivakumar and R. Wong. Linear equations in infinite matrices. Linear Algebra and its Applications, 7(1):53–62, January 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379573900360.

Tam:1975:NCP

[Tam75] Bit-Shun Tam. A note on cross-positive matrices. Linear Algebra and its Applications, 12(1):7–9, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379575901226.

Tam:1976:SRC

[Tam76] Bit-Shun Tam. Some results on cross-positive matrices. Linear Algebra and its Applications, 15(2):173–176, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900148.

Tam:1978:IS

[Tam78] Bit-Shun Shun Tam.  On indecomposable subcones.  Lin-

ear Algebra and its Applications, 22(??):33–40, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957890054X.

Tanabe:1975:NTE

[Tan75] Kunio Tanabe. Neumann-type expansions of reflexive generalized inverses of a matrix and the hyperpower iterative method. Linear Algebra and its Applications, 10(2):163–175, April 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900087.

Taussky:1972:RSM

[Tau72] Olga Taussky. The role of symmetric matrices in the study of general matrices. Linear Algebra and its Applications, 5(2):147–154, April 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379572900249.

Taussky:1975:ACR

[Tau75] Olga Taussky. Additive commutators of rational 2 × 2 matrices. Linear Algebra and its Applications, 12(1):1–6, ???? 1975.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575901214.

Truemper:1978:LUM

[TC78] K. Truemper and R. Chandrasekaran. Local unimodularity of matrixvector pairs. Linear Algebra and its Applications, 22(??):65–78, December 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379578900587.

Tchuente:1979:PCL

[Tch79] Maurice Tchuente. Parallel calculation of a linear mapping on a computer network. Linear Algebra and its Applications, 28(??):223–247, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379579901368.

Thompson:1971:ESH

[TF71] Robert C. Thompson and Linda J. Freede. On the eigenvalues of sums of Hermitian matrices. Linear Algebra and its Applications, 4(4):369–376, October 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379571900073.

Thompson:1970:PSV

[Tho70] R. C. Thompson. The l property for singular values. Linear Algebra and its Applications, 3(2):179–183, April 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379570900145.

Thompson:1971:RAS

[Tho71] R. C. Thompson. On the real and absolute singular values of a matrix product. Linear Algebra and its Applications, 4(3):243–254, July 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957190019X.

Thompson:1972:PSI

[Tho72] R. C. Thompson. Principal submatrices IX: Interlacing inequalities for singular values of submatrices. Linear Algebra and its Applications, 5(1):1–12, January 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379572900134.

Thompson:1975:DMR

[Tho75a] R. C. Thompson. Dissipative matrices and related results. Linear Algebra and its Applications, 11(2):155–169, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900555.

Thompson:1975:SVI

[Tho75b] R. C. Thompson. Singular value inequalities for matrix sums and minors. Linear Algebra and its Applications, 11(3):251–269, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900257.

Thong:1975:SDG

[Tho75c] Nguyen-Quang-Do Thong. Structure du groupe des similitudes. Linear Algebra and its Applications, 12(3):241–246, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900464.

Thomas:1976:NTP

[Tho76a] Donald H. Thomas. A natural tensor product interpolation formula and the pseudoinverse of a matrix. Linear Algebra and its Applications, 13(3):239–250, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379576900999.

Thompson:1976:BES

[Tho76b] R. C. Thompson. The behavior of eigenvalues and singular values under perturbations of restricted rank. Linear Algebra and its Applications, 13(1–2):69–78, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379576900446.

Thompson:1976:CPP

[Tho76c] R. C. Thompson. The characteristic polynomial of a principal subpencil of a Hermitian matrix pencil. Linear Algebra and its Applications, 14(2):135–177, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379576900215.

Thompson:1979:III

[Tho79a] R. C. Thompson. Interlacing inequalities for invariant factors. Linear Algebra and its Applications, 24(??):1–31, April 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901447.

Thompson:1979:PMC

[Tho79b] R. C. Thompson. Principal minors of complex symmetric and skew matrices. Linear Algebra and its Applications, 28(??):249–255, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/002437957990137X.

Thompson:1979:SVD

[Tho79c] Robert C. Thompson. Singular values and diagonal elements of complex symmetric matrices. Linear Algebra and its Applications, 26 (??):65–106, August 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379579901733.

Turkel:1977:SHD

[Tur77] E. Turkel. Symmetric hyperbolic difference schemes and matrix problems. Linear Algebra and its Applications, 16(2):109–129, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900258.

Turisco:1979:QMS

[Tur79] JoAnn Turisco. Quadratic mappings of spheres. Linear Algebra and its Applications, 23(??):261–274, February 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579901083.

Thompson:1972:RMP

[TW72] Gerald L. Thompson and Roman L. Weil. The roots of matrix pencils (Ay = λBy): existence, calculations, and relations to game theory. Linear Algebra and its Applications, 5(3):207–226, July 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379572900031.

Tromborg:1978:BDE

[TW78] B. Tromborg and S. Waldenstrøm. Bounds on the diagonal elements of a unitary matrix. Linear Algebra and its Applications, 20(3):189– 195, June 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379578900174.

Uhlig:1973:SBD

[Uhl73] Frank Uhlig. Simultaneous block diagonalization of two real symmetric matrices. Linear Algebra and its Applications, 7(4):281–289, October 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0024379573800011.

Uhlig:1974:NSF

[Uhl74] Frank Uhlig. On the nonsingular symmetric factors of a real matrix. Linear Algebra and its Applications, 8(4):351–354, August 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379574900627.

Uhlig:1976:CFP

[Uhl76] Frank Uhlig. A canonical form for a pair of real symmetric matrices that generate a nonsingular pencil. Linear Algebra and its Applications, 14(3):189–209, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379576900665.

Uhlig:1979:RTA

[Uhl79] Frank Uhlig. A recurring theorem about pairs of quadratic forms and extensions: a survey. Linear Algebra and its Applications, 25(??):219– 237, June 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/002437957990020X.

Uhrin:1977:CFC

[Uhr77] Béla Uhrin. A characterization of finite Chebyshev sequences in R^{n}. Linear Algebra and its Applications, 18(1):59–74, ????

1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900805.

Usmani:1971:PFB

[Usm71] Riaz A. Usmani. Properties of a five-band matrix and its application to boundary value problems. Linear Algebra and its Applications, 4(1):101–115, January 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379571900322.

Vandergraft:1971:GRM

[Van71] James S. Vandergraft. Generalized Rayleigh methods with applications to finding eigenvalues of large matrices. Linear Algebra and its Applications, 4(4):353–368, October 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379571900061.

Vandergraft:1976:NIL

[Van76] James S. Vandergraft. A note on irreducibility for linear operators on partially ordered finite dimensional vector spaces. Linear Algebra and its Applications, 13(1–2):139–146, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379576900525.

VandeVel:1977:NTG

[Van77] H. Van de Vel. Numerical treatment of a generalized Vandermonde system of equations. Linear Algebra and its Applications, 17(2):149– 179, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379577900350.

VanDooren:1979:CKC

[Van79] P. Van Dooren. The computation of Kronecker’s canonical form of a singular pencil. Linear Algebra and its Applications, 27(??):103–140, October 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379579900351.

Varga:1973:CBI

[Var73] Richard S. Varga. On a connection between infima of norms and eigenvalues of associated operators. Linear Algebra and its Applications, 6(??):249–256, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379573900256.

Varah:1975:LBS

[Var75] J. M. Varah. A lower bound for the smallest singular value of a matrix. Linear Algebra and its Applications, 11(1):3–5, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575901123.

Varga:1976:DDA

[Var76a] Richard S. Varga. On diagonal dominance arguments for bounding ||A^{−1}||_{∞}. Linear Algebra and its Applications, 14(3):211–217, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900677.

Varga:1976:RTD

[Var76b] Richard S. Varga. On recurring theorems on diagonal dominance. Linear Algebra and its Applications, 13(1–2):1–9, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900379.

Varga:1977:NOQ

[Var77] Richard S. Varga. A note on an open question on ω- and τmatrices. Linear Algebra and its Applications, 18(1):45–52, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900787.

vanderSluis:1979:GDP

[vdS79] A. van der Sluis. Gershgorin domains for partitioned matrices. Linear Algebra and its Applications, 26(??):265–280, August 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901812.

vanderSluis:1979:RRC

[vdSV79] A. van der Sluis and G. W. Veltkamp. Restoring rank and consistency by orthogonal projection. Linear Algebra and its Applications, 28 (??):257–278, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379579901381.

Veselic:1979:RER

[Ves79] K. Veselić. On real eigenvalues of real tridiagonal matrices. Linear Algebra and its Applications, 27(??):167–171, October 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579900399.

Vetter:1975:VSS

[Vet75] William J. Vetter. Vector structures and solutions of linear matrix equations. Linear Algebra and its Applications, 10(2):181–188, April 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900105.

Vidav:1976:GIS

[Vid76] Ivan Vidav. The group of isometries and the structure of a finitedimensional Banach space. Linear Algebra and its Applications, 14(3):227–236, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379576900690.

Vitoria:1979:MND

[Vit79] José Vitória. Matricial norms and the differences between the zeros of determinants with polynomial elements. Linear Algebra and its Applications, 28(??):279–283, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579901393.

vKempen:1970:VES

[vK70] H. P. M. v. Kempen. Variation of the eigenvalues of a special class of Hermitian matrices upon variation of some of its elements. Linear Algebra and its Applications, 3(2):263–273, April 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379570900194.

vanLeeuwen:1978:SEP

[vLvEB78] J. van Leeuwen and P. van Emde Boas. Some elementary proofs of lower bounds in complexity theory. Linear Algebra and its Applications, 19(1):63–80, ???? 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379578900058.

Wall:1975:GIS

[Wal75] J. R. Wall. Generalized inverses of stochastic matrices. Linear Algebra and its Applications, 10(2):147–154, April 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379575900063.

Wang:1979:GC

[Wan79] Kai Wang. On the generalizations of circulants. Linear Algebra and its Applications, 25(??):197–218, June 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579900193.

Wasow:1977:ACM

[Was77] Wolfgang Wasow. Arnold’s canonical matrices and the asymptotic simplification of ordinary differential equations. Linear Algebra and its Applications, 18(2):163–170, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379577900714.

Watford:1972:SCG

[Wat72] Luck J. Watford, Jr. The Schur complement of a generalized Mmatrix. Linear Algebra and its Applications, 5(3):247–255, July 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379572900067.

Watkins:1976:LMP

[Wat76] William Watkins. Linear maps that preserve commuting pairs of matrices. Linear Algebra and its Applications, 14(1):29–35, ???? 1976. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379576900604.

Webber:1974:MSD

[Web74] Robert P. Webber. Matrix semigroups determined by unitarily invariant norms. Linear Algebra and its Applications, 8(2):141–156, April 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379574900524.

Wehrfritz:1977:FCH

[Weh77] B. A. F. Wehrfritz. Finite central height in linear groups. Linear Algebra and its Applications, 17(1):59–64, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379577900416.

Wernick:1970:PGE

[Wer70] Robert J. Wernick. Parametrized games and the eigenproblem. Linear Algebra and its Applications, 3(3):311–346, July 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379570900030.

Werner:1973:EUP

[Wer73] B. Werner. Einschließung des positiven eigenvektors einer nichtnegativen, irreduziblen matrix. (German) [Enclosure of positive eigenvector of a nonnegative irreducible matrix]. Linear Algebra and its Applications, 7(4):351–359, October 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0024379573800060.

Werner:1979:MMG

[Wer79] Hans-Joachim Werner. On the matrix monotonicity of generalized inversion. Linear Algebra and its Applications, 27(??):141–145, October 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579900363.

Westwick:1972:SLT

[Wes72] R. Westwick. Spaces of linear transformations of equal rank. Linear Algebra and its Applications, 5(1):49–64, January 1972. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379572900183.

Wibker:1979:ESR

[WHG79] E. A. Wibker, R. B. Howe, and J. D. Gilbert. Explicit solutions to the reverse order law (AB)^{+} = B_{mr}^{−}A_{lr}^{−}. Linear Algebra and its Applications, 25(??):107–114, June 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379579900107.

Wiitala:1978:FIC

[Wii78] Stephen A. Wiitala. Factorization of involutions in characteristic two orthogonal groups: an application of the Jordan form to group theory. Linear Algebra and its Applications, 21(1):59–64, July 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437958790200X.

Williams:1970:BRL

[Wil70] A. C. Williams. Boundedness relations for linear constraint sets. Linear Algebra and its Applications, 3(2):129–141, April 1970. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379570900091.

Wilkinson:1975:AHA

[Wil75a] J. H. Wilkinson. Alston S. Householder award. Linear Algebra and its Applications, 11(1):1, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379575901111.

Wilson:1975:BIN

[Wil75b] Peter B. Wilson. Bounds for iterated normal matrices. Linear Algebra and its Applications, 10(1):7–18, February 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379575900907.

Willoughby:1977:IMP

[Wil77] R. A. Willoughby. The inverse M-matrix problem. Linear Algebra and its Applications, 18(1):75–94, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379577900817.

Wilkinson:1979:KCF

[Wil79] J. H. Wilkinson. Kronecker’s canonical form and the QZ algorithm. Linear Algebra and its Applications, 28(??):285–303, December 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/002437957990140X.

Wimmer:1974:ITM

[Wim74] Harald K. Wimmer. Inertia theorems for matrices, controllability, and linear vibrations. Linear Algebra and its Applications, 8(4):337–343, August 1974. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0024379574900603.

Wimmer:1975:GTL

[Wim75] Harald K. Wimmer. Generalizations of theorems of Lyapunov and Stein. Linear Algebra and its Applications, 10(2):139–146, April 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379575900051.

Wimmer:1978:JKR

[Wim78] Harald K. Wimmer. Jordan–Ketten und realisierungen rationaler matrizen. Linear Algebra and its Applications, 20(2):101–110, ???? 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379578900447.

Winograd:1971:MM

[Win71a] S. Winograd. On multiplication of 2×2 matrices. Linear Algebra and its Applications, 4(4):381–388, October 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379571900097.

Winograd:1971:ACI

[Win71b] S. Winograd. On the algebraic complexity of inner product. Linear Algebra and its Applications, 4(4):377–379, October 1971. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379571900085.

Kuo:1977:MPI

[wK77a] I wen Kuo. The Moore–Penrose inverses of singular M-matrices. Linear Algebra and its Applications, 17(1):1–14, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900386.

Kuo:1977:NFS

[wK77b] I wen Kuo. A note on factorizations of singular M-matrices. Linear Algebra and its Applications, 16(3):217–220, ???? 1977.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379577900040.

Yarlagadda:1978:SM

[Yar78] R. Yarlagadda. Stabilization of matrices. Linear Algebra and its Applications, 21(3):271–288, September 1978. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0024379578900885.

Yamamoto:1979:IBM

[YI79] Tetsuro Yamamoto and Yasuhiko Ikebe. Inversion of band matrices. Linear Algebra and its Applications, 24(??):105–111, April 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901514.

Young:1979:NPM

[You79] N. J. Young. Norms of powers of matrices with constrained spectra. Linear Algebra and its Applications, 23(??):227–244, February 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901058.

Zellini:1979:SPC

[Zel79a] Paolo Zellini. On some properties of circulant matrices. Linear Algebra and its Applications, 26(??):31–43, August 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379579901708.

Zellini:1979:OCS

[Zel79b] Paolo Zellini. On the optimal computation of a set of symmetric and persymmetric bilinear forms. Linear Algebra and its Applications, 23 (??):101–119, February 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/002437957990096X.

Zenger:1977:MSI

[Zen77] Christoph Zenger. Minimal subadditive inclusion domains for the eigenvalues of matrices. Linear Algebra and its Applications, 17 (3):233–268, ???? 1977. CODEN LAAPAW. ISSN 0024-3795 REFERENCES 130

(print), 1873-1856 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0024379577900635.